### v1

# Online Resource Allocation via Static Bundle Pricing

Abstract
Online Resource Allocation addresses the problem of efficiently allocating limited resources to buyers with incomplete knowledge of future requests. In our setting, buyers arrive sequentially demanding a set of items, each with a value drawn from a known distribution. We study environments where buyers’ valuations exhibit complementarities. In such settings, standard item-pricing mechanisms fail to leverage item multiplicities, while existing static bundle-pricing mechanisms rely on problem-specific arguments that do not generalize.
We develop a unified technique for online resource allocation with complementarities for three domains: (i) single-minded combinatorial auctions with maximum bundle size $d$, (ii) general single-minded combinatorial auctions, and (iii) a graph-based routing model in which buyers request to route a unit of flow from a source node $s$ to a target node $t$ in a capacitated graph. Our approach yields static and anonymous bundle-pricing mechanisms whose performance improves exponentially with item capacities. For the $d$-single-minded setting with minimum item capacity $B$, we obtain an $O(d^{1/B})$-competitive mechanism, recovering the known $O(d)$ bound for unit capacities ($B=1$) and achieving exponentially better guarantees as capacities grow. For general single-minded combinatorial auctions and the graph-routing model, we obtain $O(m^{1/(B+1)})$-competitive mechanisms, where $m$ is the number of items.
We complement these results with information-theoretic lower bounds. We show that no online algorithm can achieve a competitive ratio better than $\Omega((m/\ln m)^{1/(B+2)})$ in the general single-minded setting and $\Omega((d/\ln d)^{1/(B+1)})$ in the $d$-single-minded setting. In doing so, we reveal a deep connection to the classical extremal combinatorics problem of determining the maximum number of qualitatively independent partitions of a ground set.
footnotetext: This work has been partially supported by project MIS 5154714 of the National Recovery and Resilience Plan Greece 2.0, funded by the European Union under the NextGenerationEU Program.
1 Introduction
Online Resource Allocation is a central problem in Economics and Operations Research. A seller faces a sequence of buyers, each demanding a set of resources. The seller must make irrevocable allocation decisions with limited information about future arrivals. In this work, we study the setting in which the seller has access to the distribution of buyers’ valuations, placing the problem within the prophet-inequality framework.
The study of prophet inequalities originates in the classical single-item setting. Krengel, Sucheston, and Garling [krengel1977] and independently Samuel-Cahn [samuel1984comparison] established the foundational results comparing online stopping rules to a prophet who observes all realizations in advance. Hajiaghayi et al. [haji1007] later connected prophet inequalities to mechanism design, showing how prophet-style guarantees can be leveraged to design truthful posted-price mechanisms. Building on this connection, Chawla et al. [chawla2009] developed a theory of sequential posted pricing mechanisms.
Following these foundational results, research on prophet inequalities expanded rapidly, branching into richer allocation environments and more complex valuation models. In particular, combinatorial auctions emerged as a central and intensively studied setting within the prophet-inequality framework. For submodular and XOS valuations, Feldman et al. [fgl-balanced14] gave a $2$-competitive posted-price mechanism. For subadditive valuations, Dütting et al. [duttingloglogm] obtained, via item-pricing, an $O(\log\log m)$-competitive mechanism (where $m$ is the number of items), a guarantee that was recently improved by Correa and Cristi [Correa2023ACF], who achieved a $6$-competitive prophet inequality. For MPH-$k$ valuations, Braun and Kesselheim [braun2022simplifiedprophetinequalitiescombinatorial] designed an item-pricing mechanism with competitive ratio $2k+\sqrt{2k(k-1)}-1$.
In this work, we focus on single-minded combinatorial auctions, a setting that captures strong complementarities: each buyer derives value only if they obtain their entire desired bundle. Single-minded combinatorial auctions capture the most extreme form of complementarities and arise naturally in applications where partial allocation provides no utility. Their simplicity makes them a fundamental testbed for understanding how complementarities interact with online decision-making and posted-price mechanisms. A canonical example is pricing goods in a grocery store. A shopper attempting to bake a cake gains value only if they obtain all necessary ingredients; eggs and milk provide zero utility if flour is missing.
Extending the grocery store analogy, many real systems feature multiple copies of each item rather than unique goods. A growing line of work examines how such multiplicities affect the performance of pricing mechanisms. Alaei [multi-alei] showed that, in the single-item setting with capacity $k$, one can achieve $1-O(1/\sqrt{k})$ competitiveness. More recent results by Jiang et al. [multi-jiang] and Chawla et al. [multi-suchi] give matching upper and lower bounds for all values of $k$, while Chawla et al. [chawla25] continued this line of work for XOS valuations. Recently, Ghuge et al. [ghuge2025] studied online resource allocation with large capacities under stochastic arrivals using dynamic pricing mechanisms. They show that near-optimal welfare can be achieved using only a single sample from each distribution, highlighting that high capacities substantially mitigate uncertainty even with limited information.
In contrast, our work focuses on static mechanisms, in which the seller commits to a menu of prices before any buyer arrives and does not adjust them over time. Although such mechanisms are simple, transparent, and widely deployed in practice, they face significant limitations in the presence of complementarities. The seminal work of Feldman et al. [fgl-balanced14] established strong impossibility results for item pricing in combinatorial auctions: when buyers’ valuations require complements, no item-pricing mechanism can be less than $\Omega(m)$-competitive. When parameterizing by the maximum bundle size $d$, Correa et al. [correa-d] obtained a tight $(d+1)$-competitive guarantee.
Motivated by these challenges, Chawla et al. [chawla18] initiated the study of bundle pricing in settings with complementarities, first in the context of online allocation of intervals and later for paths in trees. Their results demonstrated that richer pricing menus can circumvent the barriers inherent to item pricing. Moreover, their work highlighted that static pricing mechanisms admit progressively stronger guarantees as the number of available copies of each resource increases.
This raises two fundamental questions:
What static pricing mechanisms are appropriate for environments with complementarities?
How does the availability of multiple item copies affect the achievable performance of such mechanisms?
In this work, we develop a unified approach for prophet inequalities in the presence of complementarities. Our mechanisms are static (the corresponding menus are fully determined before any arrivals occur) and anonymous (buyers face the same menu). We show that this approach is applicable to the following domains:
  1. 1.
    the $d$-single-minded setting, where each buyer demands a bundle of size at most $d$
  2. 2.
    general single-minded combinatorial auctions, with arbitrary bundle sizes
  3. 3.
    the graph-routing setting, where buyers seek any feasible $s$-$t$ path in a capacitated graph.
Across all of these environments, a pattern emerges: bundle prices become exponentially more powerful as item capacities increase. Even small increases in capacity can dramatically improve achievable competitive ratios, revealing a sharp structural transition in the difficulty of online allocation with complementarities.
At the same time, we complement our positive results with information-theoretic lower bounds for both the $d$-single-minded and the general single-minded settings. Our lower-bound construction is obtained by embedding the classical hard instance of Babaioff et al. [babaioff] into the single-minded combinatorial auction framework. This embedding reveals an intriguing connection between our impossibility results and an extremal combinatorics problem: determining the maximum number of pairwise qualitatively independent partitions of a ground set, a question first studied by Rényi [renyi]. Together, these lower bounds show that the dependence of our guarantees on the number of items $m$ and the minimum multiplicity $B$ is essentially unavoidable.
1.1 Technical Overview and Contributions
We study online allocation problems with strong complementarities under distributional information, focusing on static and anonymous bundle-pricing mechanisms. All guarantees hold in expectation over both the randomness of the mechanism and buyers’ realized values, and the arrival order is chosen by an almighty adversary who observes the entire realization and the mechanism’s internal randomness.
Our starting point is the ex-ante linear relaxation, which upper bounds the expected welfare of the prophet, i.e. the expected value over the realizations welfare of the optimal allocation. Before solving this LP, we scale down all item capacities by a factor $\gamma$, creating a slack that later absorbs the fluctuations created by the randomness of the realization. This scaled relaxation guides the design of our pricing mechanism.
A notable feature of the optimal fractional solution is its simple and rigid structure. For any buyer, the LP allocates probability mass to values in an almost threshold-like form: below some cutoff the allocation is zero, above it the LP fully saturates the value distribution, and at most one crucial value may receive partial mass. When aggregated over all buyers who demand the same bundle $S$, these patterns define a canonical price level $w_{S}$ that dictates how the menu treats $S$.
This structure directly informs our menu construction. The mechanism posts a static, anonymous menu containing multiple copies of each bundle $S$, priced either at the crucial value $w_{S}$ or at the next higher value in the buyers’ value support, $w_{S}+1$. The guiding objective is to emulate the fractional solution as closely as possible: ideally, a buyer $b$ with realized value $v$ would find an available copy of their desired bundle with probability $x_{b,v}/q_{b,v}$, independently of other buyers, where $x_{b,v}$ is the LP mass assigned to the buyer-value pair $(b,v)$ and $q_{b,v}=\mathbb{P}\!\left[v_{b}=v\right]$.
Such independent acceptance probabilities, however, cannot be implemented directly using a static and anonymous menu. Instead, the mechanism must commit ex-ante to a fixed number of copies of each bundle at each price. Guided by the LP solution, we post no copies of bundle $S$ at prices below $w_{S}$, and essentially unlimited copies at prices strictly above $w_{S}$, relying on the scarcity of high-value buyers to limit congestion. The number of copies offered exactly at price $w_{S}$ is more delicate. Offering too many copies risks exhausting item capacities and blocking high-value buyers later in the sequence, while offering too few results in a significant loss relative to the fractional benchmark. To balance these effects, we randomize between offering $\lfloor x_{S,w_{S}}\rfloor$ and $\lceil x_{S,w_{S}}\rceil$ copies of bundle $S$ at price $w_{S}$, where $x_{S,w_{S}}$ denotes the total LP mass assigned to buyers demanding $S$ with value $w_{S}$.
A desirable property of our menu is price subadditivity: for every bundle $S$, the posted price of $S$ should be no greater than the total price of any collection of bundles whose union contains $S$. This guarantees that a buyer who desires $S$ always prefers purchasing a copy of $S$ rather than assembling a larger covering collection. Enforcing price subadditivity, however, is nontrivial. The prices suggested by the LP solution need not satisfy this property, and the buyers’ original value supports may not contain appropriate price levels that would permit a subadditive menu without sacrificing performance. To address this issue, we first augment each buyer’s value support with carefully chosen auxiliary price points. We then apply a post-processing step to the LP solution that enforces subadditivity before executing the menu-construction procedure described above.
To analyze the mechanism, we introduce an unconstrained algorithm: a hypothetical process that accepts every buyer who selects an available menu entry, ignoring item-capacity constraints. This auxiliary algorithm is easier to analyze and satisfies two key properties:
  1. 1.
    Its expected welfare is a constant fraction of the scaled LP optimum, denoted by $\textnormal{{FracOpt}}_{\gamma}$ throughout the paper.
  2. 2.
    For every item $e$, the expected number of sold copies that use $e$ is at most $c(e)/\gamma$.
The welfare of the actual mechanism equals the welfare of the unconstrained algorithm minus the value of blocked buyers-those accepted by the unconstrained algorithm but rejected when enforcing capacity constraints. Property (1) implies that it suffices to bound the expected welfare lost due to blocking as a constant fraction of the unconstrained value, while Property (2) allows us to control the expected contribution of blocked buyers.
d-Single-Minded Setting.
In the $d$-single-minded setting, there is a set $\mathcal{M}$ of $m$ items, each with multiplicity specified by a capacity function $c:\mathcal{M}\to\mathbb{Z}_{>0}$. Buyers arrive online, each drawing a value $v_{b}\sim\mathcal{D}_{b}$, and derive utility $v_{b}$ if and only if they receive their entire desired bundle $S_{b}$, where $|S_{b}|\leq d$. Both the bundle $S_{b}$ and the distribution $\mathcal{D}_{b}$ are known in advance.
In this setting, we scale the item capacities in the LP relaxation by a factor $\gamma=e(10d)^{1/B}.$ Using a concentration argument, we show that for any fixed item, the probability that it becomes congested is at most $1/(10d)$. Taking a union bound over the at most $d$ items contained in a buyer’s desired bundle, we obtain that any bundle allocated by the unconstrained algorithm is congested with probability at most $0.1$. We leverage this bound to show that the expected value lost due to blocked buyers is at most a $1/10$ fraction of the value of the scaled fractional optimum. Combined with the properties of the unconstrained algorithm, this yields the following guarantee.
Theorem 1.1.
For the $d$-single-minded case with minimum item multiplicity $B$, our mechanism achieves an $O(d^{1/B})$-competitive static and anonymous posted-price mechanism.
General Single-Minded Setting.
We next consider the fully general single-minded setting, in which buyers may demand arbitrary bundles of items. Relative to the $d$-single-minded case, the absence of an a priori bound on bundle size requires more aggressive capacity scaling. Specifically, we scale down item capacities in the LP relaxation by a factor $\gamma=e(20m)^{1/(B+1)}.$ The slack created by this scaling allows us to show that the expected cardinality of the set of blocked buyers is at most $1/20$.
A further complication arises from the possibility of unbounded valuations: even a single blocked buyer with a very large value could dominate the welfare loss, rendering bounds on the number of blocked buyers insufficient. To address this issue, we augment the main mechanism with a simple small-market component that, with a carefully chosen probability, posts a single bundle at an extremely high price. This auxiliary mechanism captures a constant fraction of the welfare contributed by very large values, ensuring that heavy-tailed realizations cannot overwhelm the performance guarantee of the LP-guided menu. The main mechanism handles the structured, high-probability portion of the market, while the small-market mechanism isolates and stabilizes the contribution of extreme values. Together, the combined mechanism yields the following guarantee.
Theorem 1.2.
For general single-minded combinatorial auctions with minimum multiplicity $B$, our mechanism achieves an $O\!\left(m^{1/(B+1)}\right)$-competitive static and anonymous posted-price mechanism.
Graph Setting.
Finally, we extend our framework to the network-routing domain, in which each buyer seeks to route a unit of flow from a source node $s$ to a target node $t$ in a capacitated graph $G=(V,E)$. We view this setting as a natural generalization of the general single-minded case: rather than demanding a single bundle, each buyer can be satisfied by any feasible $s$-$t$ path. The challenge arising by the extension is that the mechanism must select a specific routing for each buyer.
Once again, the structure of the LP relaxation plays a central role. We conceptually treat each source-target pair $(s,t)$ as an abstract bundle and group buyers according to their requested pair. Applying the same menu-construction procedure as in the general single-minded setting determines how many copies are posted to serve each $(s,t)$ pair, thereby inheriting the welfare guarantees of the unconstrained algorithm.
It remains to assign an explicit $s$-$t$ path to each such copy. To this end, we independently sample a path for each copy according to a carefully chosen distribution derived from the LP solution. Using a roulette-style sampling argument, we show that this randomized routing preserves the expected load guarantees on every edge, matching those obtained in the single-minded analysis. This yields the following result.
Theorem 1.3.
For the graph-routing setting with minimum edge capacity $B$, our mechanism achieves an $O\!\left(m^{1/(B+1)}\right)$-competitive static posted-price mechanism, where $m=|E|$.
Lower bounds.
Our impossibility results follow a standard lower-bound template for prophet inequalities (see, e.g., [chawla2009, babaioff]). In this template, buyers are partitioned into many groups of equal size $t$. Buyers arrive sequentially, each has value in $\{0,1\}$ with $\mathbb{P}\!\left[v=1\right]=1/t$, and feasibility is engineered so that we can select buyers from at most one group. Since the expected number of buyers realized in a fixed group is $1$, after committing to a group, the online algorithm can extract only constant expected value from that group and thus gains at most $O(1)$ value. In contrast, if we create about $t^{t}$ groups, then with constant probability there exists a fully active group (all $t$ buyers have value $1$), allowing the prophet to obtain welfare on the order of $t$. The main challenge is to embed this template into single-minded combinatorial auctions, where feasibility is induced by item capacities.
For unit capacities, we associate each group with a partition of the $m$ items into $t$ disjoint classes. The $t$ buyers in the group are single-minded, with buyer $i$ demanding exactly the $i$-th class. Disjointness within a partition guarantees that within a group all $t$ buyers can simultaneously be served. To prevent serving buyers from two different groups, we choose a family of partitions that are pairwise qualitatively independent (QI), a notion introduced by Rényi [renyi], where every class of one partition intersects every class of another. Then any allocation that serves one buyer from each of two groups must allocate some common item twice, contradicting unit capacity. This realizes the "at most one group" constraint of the template using item feasibility.
With $B$ copies per item, we generalize "at most one group" to "at most $B$ groups". The right combinatorial notion is (B+1)(B\!+\!1)-way qualitative independence: for any B+1B\!+\!1 partitions and any choice of one class from each, their intersection is nonempty. Using such a family, any allocation that serves buyers from B+1B\!+\!1 distinct groups must allocate some item to B+1B\!+\!1 buyers, contradicting the $B$-copy constraint; meanwhile, any fixed set of $B$ groups is simultaneously feasible because each item appears in at most one demanded class per group.
Poljak and Tuza [poljak] give the following lower bound on the number $r$-way qualitatively independent partitions of $m$ items into $t$ classes (denoted by $g(m,t,r)$):
$g(m,t,r)\geq\frac{er}{t}\cdot\exp\left(\frac{m}{rt^{r}}\right)$
thus, choosing
$t\;=\;\left\lfloor\left(\frac{m}{2B\ln m}\right)^{1/(B+2)}\right\rfloor,$
via simple algebra we can prove that there exists a (B+1)(B\!+\!1)-way QI family of size at least $Bt^{t}$, thus we can instantiate $N=Bt^{t}$ groups. In our instance, each group is fully active with probability $\left(1/t\right)^{t}$ and if we let $X$ be the number of fully active groups we get that $\mathbb{E}\left[X\right]=N\cdot\left(1/t\right)^{t}=B$, and with constant probability it holds $X\geq B$. The prophet then serves all $tB$ active buyers, while any online algorithm is restricted to at most $B$ groups and extracts only constant expected value per served group.
Theorem 1.4.
For the general single-minded setting with $m$ items and minimum item capacity $B\geq 1$, every online algorithm has competitive ratio,
$\Omega\left(\left(\frac{m}{\ln m}\right)^{1/(B+2)}\right).$
To obtain a lower bound in the $d$-single-minded setting, we additionally require every demanded bundle to have size at most $d$. This leads us to introduce the notion of balanced QI partitions, in which every class has size $\leq d$. We show that the probabilistic construction of [poljak] extends to this balanced setting (sampling uniformly from balanced labelings), yielding large (B+1)(B\!+\!1)-way QI families with all classes of size about $m/t$. Setting $m=dt$ and choosing
$t\;=\;\left\lfloor\left(\frac{d}{2B\ln d}\right)^{1/(B+1)}\right\rfloor,$
the same group-commitment argument goes through.
Theorem 1.5.
For the $d$-single-minded setting with maximum desired bundle size $d$ and minimum item capacity $B\geq 1$, every online algorithm has competitive ratio,
$\Omega\left(\left(\frac{d}{\ln d}\right)^{1/(B+1)}\right).$
For unit capacities, there exist stronger combinatorial constructions that improve the dependence on $m$ beyond what the general probabilistic bound yields. Using the explicit construction of [poljak], we build $t^{t}$ pairwise QI $t$-partitions on $m=p^{3}$ items with $t=p$, for some sufficiently large prime $p$, eliminating the $\ln m$ loss.
Theorem 1.6.
For the single-minded setting with $m$ items and unit capacities, every online algorithm has competitive ratio,
$\Omega\left(m^{1/3}\right).$
Finally, we relate pairwise qualitative independence to the Alon-Alweiss measure (AAM) introduced by [AA20, SVW22], which yields even larger QI families on slightly fewer items and thus a mild exponent improvement.
Theorem 1.7.
For the single-minded setting with $m$ items and unit capacities, every online algorithm has competitive ratio,
$\Omega\left(m^{\,1/3+c/\log\log m}\right)$
for some absolute constant $c>0$.
1.2 Further Related Work
Online Contention Resolution Schemes.
We work with the ex-ante linear programming relaxation and interpret the fractional value assigned to each buyer-value pair as a target acceptance probability. The online realizations then play the role of a randomized rounding step. Our mechanism is responsible for enforcing feasibility ex-post, while preserving as much of the fractional optimum as possible.
This perspective is closely related to the framework of contention resolution schemes (CRSs), introduced by Chekuri, Vondrák, and Zenklusen for submodular maximization under combinatorial constraints [crs-chekuri]. In this framework, a CRS is given a fractional point $x^{*}$ in the relaxation polytope and samples a random set $R$ according to the product distribution with marginals $x^{*}$. Since the resulting set $R$ may violate feasibility constraints, the scheme subsequently discards selected elements to obtain a feasible subset $R^{\prime}\subseteq R$. Later, Feldman, Svensson, and Zenklusen [ocrs] introduced the online counterpart of this framework, termed online contention resolution schemes (OCRSs), which apply to a wide range of online settings, including prophet inequalities. In their work, they showed that the existence of a $c$-balanced OCRS implies a $c$-competitive prophet inequality. More recently, Lee and Singla [optimal-ocrs-singla] showed that a $c$-competitive prophet inequality also yields a $c$-competitive ex-ante OCRS.
Static versus Dynamic Pricing.
A related line of work is due to Rubinstein [rubinstein-beyond-matroids], who studied prophet inequalities under general downward-closed feasibility constraints. He obtained an $O(\log n\cdot\log r)$-competitive algorithm, where $n$ denotes the number of arriving elements (corresponding to buyers in our setting) and $r$ the maximum cardinality of any feasible solution. Rubinstein’s algorithm observes each realized value upon arrival and accepts the element if doing so does not decrease a carefully designed dynamic potential function. Since this potential function depends on the current accepted set, the resulting acceptance rule is inherently dynamic. Recently, Banihashem et al. [krysta] showed that any prophet inequality has an implementation using a (sequential) posted price mechanism.
Lower Bounds.
The general $\Omega(\log n/\log\log n)$ lower bound in prophet inequalities with $n$ buyers that is used as a template for our main lower bound results has been introduced in other forms by Babaioff, Immorlica, and Kleinberg [babaioff] and by Chawla et al. [chawla2009]. For the $d$-single-minded cases, lower bounds for prophet inequalities under the intersection of $q$ partition matroids also imply hardness for $q$-single-minded prophet inequality instances: Kleinberg and Weinberg [klein-wein12] gave a $\Omega(\sqrt{q})$ lower bound and Saxena, Velusamy, and Weinberg [SVW22] strengthened this to $q^{1/2+\Omega(1/\log\log q)}$ using techniques of Alon and Alweiss [AA20].
2 Preliminaries
Model.
The input consists of a set of $m$ distinct items $\mathcal{M}$ and copies per item specified by the function $c:\mathcal{M}\to\mathbb{Z}_{>0}$. Let $B=\min_{e\in\mathcal{M}}c(e)$ denote the minimum number of copies among all items $e\in\mathcal{M}$. We are also given a set $\mathcal{N}$ of $n$ buyers. We index buyers by $b\in\{1,\ldots,n\}$. Each buyer $b$ is associated with a fixed set of items $S_{b}$ known to the seller. In addition, each buyer $b$ is associated with a valuation function $v^{(b)}:2^{\mathcal{M}}\to\mathbb{R}_{\geq 0}$ such that $v^{(b)}(S)=v_{b}$ if $S_{b}\subseteq S$ and $v^{(b)}(S)=0$ otherwise. Therefore, buyers are single-minded and thus fully described by the value $v_{b}$ and set $S_{b}$.
For the rest of the paper we will consider the Bayesian setting where $v_{b}$ is drawn from a distribution $\mathcal{D}_{b}$ that is known to the seller and independent across buyers. We will assume that the distributions are discrete with finite support $\mathcal{S}=\{0,\ldots,R\}$ with $R\in\mathbb{Z}_{>0}$, thus for each buyer $b$ and each $v\in\mathcal{S}$ we have $q_{b,v}>0$. In Appendix A we remove this assumption and we allow buyers to have different supports. Buyers arrive sequentially, each revealing their realized value upon arrival. We assume that the order of arrival is determined by an almighty adversary. Specifically, the almighty adversary can choose the arrival order after observing all realizations of the valuations and the random choices of the algorithm. Our objective is to maximize the social welfare, i.e., the sum of the values derived from the sets of items allocated to the buyers.
The class of mechanisms we consider is static and anonymous bundle pricing mechanisms. More specifically, before the arrival of the buyers, we offer a menu $M$ that contains multiple entries of the form $(T_{j},p_{j},c_{j})$, with $T_{j}\in 2^{\mathcal{M}}$, $p_{j}\in\mathbb{R}_{\geq 0}$ and $c_{j}\in\mathbb{Z}_{>0}$. In other words, we offer $c_{j}$ copies of the bundle of items $T_{j}$ at price $p_{j}$ each. Upon arrival, each buyer $b$ chooses a subset $J$ of available menu entries that maximizes their utility, where their utility is $v_{b}$ if $S_{b}\subseteq\bigcup_{j\in J}T_{j}$ and 0 otherwise, minus the total payment $\sum_{j\in J}p_{j}$. Note that a menu entry $(T_{j},p_{j},c_{j})$ can be chosen only if fewer than $c_{j}$ copies of this entry have been sold so far and copies of each item in $T_{j}$ are still available. The above class of mechanisms is static since it is determined upfront (before the realization of the buyers’ values) and does not change during the run of the algorithm. Also, it is anonymous because we offer the same menu to each buyer.
Notation.
We use the notation $x]\coloneqq\{1,\ldots,x\}$ for any $x\in\mathbb{Z}_{>0}$. For $v\in\mathcal{S}$ define $q_{b,v}\coloneqq\mathbb{P}_{v_{b}\sim\mathcal{D}_{b}}\!\left[v_{b}=v\right]$ and let $\mathcal{D}\coloneq\prod_{b\in[n]}\mathcal{D}_{b}=\mathcal{D}_{1}\times\ldots\times\mathcal{D}_{n}$. Let $I\sim\mathcal{D}$ be an instance of buyers’ values. For each buyer $b$, we denote by $v_{b}$ the random variable corresponding to its value. Note that, throughout the paper, whenever an index of the variables $x_{b,v}$ where $b\in[n],\ v\in\mathcal{S}$ is omitted, we implicitly assume summation over that index (unless stated otherwise). For example, $x_{b}=\sum_{v\in\mathcal{S}}x_{b,v}$ and $x_{v}=\sum_{b\in[n]}x_{b,v}$. For an event $\mathcal{E}$ we define $\mathbbm{1}{\left\{\mathcal{E}\right\}}$ as its indicator function. Finally, let $\ell_{A}(e)$ denote the load of item $e$ due to buyers in $A$, i.e. $\ell_{A}(e)=\sum_{b\in A}\mathbbm{1}{\left\{e\in S_{b}\right\}}$.
2.1 Linear Program
Like in many other works in the stochastic setting, we will compare against the expected offline optimal social welfare Opt through a fractional relaxation of the problem. Consider the following LP relaxation:
$\max\quad$ $\sum_{b\in[n]}\sum_{v\in\mathcal{S}}x_{b,v}\cdot v$
s.t. $\sum_{b\in[n]:e\in S_{b}}\sum_{v\in\mathcal{S}}x_{b,v}\leq c(e),\qquad\forall e\in\mathcal{M}$
$0\leq x_{b,v}\leq q_{b,v},\qquad\forall b\in[n],\ \forall v\in\mathcal{S}$
Note that in the above LP, the capacity constraints are enforced only in expectation (ex-ante), so $x$ represents a feasible ex-ante fractional allocation. Let FracOpt be the optimal value of the previous LP. Generally, fractional relaxations upper bound the optimal welfare Opt, (see for instance Lemma 2.1 in [chawla18]). For completeness, in Appendix G we prove the following.
Lemma 2.1.
$\textnormal{{FracOpt}}\geq\textnormal{{Opt}}$.
2.2 Structure of the LP solution
Before analyzing our bundle pricing, we first establish several structural properties of the fractional optimal solution of the LP. The key conceptual tool for understanding this structure is the notion of value exchange.
Suppose for the moment that we ignore the constraints $x_{b,v}\leq q_{b,v}$. Consider two variables $x_{b^{\prime},v^{\prime}}>0$ and $x_{b^{\prime\prime},v^{\prime\prime}}$ with $S_{b^{\prime\prime}}\subseteq S_{b^{\prime}}$. If we decrease $x_{b^{\prime},v^{\prime}}$ by a small amount (mass) $\delta>0$ and increase $x_{b^{\prime\prime},v^{\prime\prime}}$ by the same amount, then the capacity constraints remain satisfied, since allocating mass on a subset of a set consumes no more capacity than allocating the same mass on the original set. This exchange strictly increases the LP objective whenever $v^{\prime\prime}>v^{\prime}$, as we shift mass from a lower-value variable to a higher-value one without affecting feasibility. Thus, in an optimal LP solution $x$, no such improving exchange should be possible. The only way such an exchange can be blocked is by the constraints $x_{b^{\prime\prime},v^{\prime\prime}}\leq q_{b^{\prime\prime},v^{\prime\prime}}$. Consequently, for two variables $0<x_{b^{\prime},v^{\prime}}$ and $x_{b^{\prime\prime},v^{\prime\prime}}$ with $S_{b^{\prime\prime}}\subseteq S_{b^{\prime}}$, whenever $v^{\prime\prime}>v^{\prime}$, we must have $x_{b^{\prime\prime},v^{\prime\prime}}=q_{b^{\prime\prime},v^{\prime\prime}}$. In the following section, we formalize this intuition, first for each buyer individually and then for each set $S$.
Per Buyer.
Intuitively, our goal is to determine a price to offer each buyer. If we did not require the mechanism to be anonymous or static, a natural approach would be the following: upon the arrival of buyer $b$ with realized value $v$, we accept them with probability $x_{b,v}/q_{b,v}$. Recall that $x_{b,v}$ can be interpreted as the ex ante probability mass of allocating $S_{b}$ to buyer $b$ when $v_{b}=v$. Thus this acceptance rule ensures that buyer $b$ is served with unconditional probability $q_{b,v}\cdot(x_{b,v}/q_{b,v})=x_{b,v}$, matching the LP’s marginal allocation in expectation. Of course, this does not yield a feasible mechanism, because the LP capacity constraints hold only in expectation and may be violated ex-post. The subsequent analysis will show how to overcome this limitation.
From this perspective, the LP solution guides our pricing decisions. Interpreting acceptance probabilities as price thresholds, the LP suggests which values should be priced out and which should be encouraged. If the LP assigns zero weight to a particular value, we should not accept buyers at that level, suggesting a higher price in a static-price mechanism. Conversely, if the LP assigns the maximum feasible weight to a value, this suggests that the corresponding price should not be set above that level. However, there may exist some intermediate crucial values where neither condition applies. Properly handling these crucial values is essential to the design of our mechanism.
Formally, fix a buyer $b$ and consider the vector $X^{(b)}=[x_{b,v}]_{v\in\mathcal{S}}$, whose entries correspond to the optimal fractional LP allocation for buyer $b$, ordered in increasing order of $v$.
Definition 2.2.
A value $v\in\mathcal{S}$ is tight for buyer $b$ if $x_{b,v}=q_{b,v}>0$, and crucial if $0<x_{b,v}<q_{b,v}$.
The following lemma, proved in Appendix G, characterizes the structure of the vector $X^{(b)}$. The proof relies on the notion of value exchange introduced above.
Lemma 2.3.
For each buyer $b$, the vector $X^{(b)}$ consists of three consecutive blocks:
  • a (possibly empty) prefix of values $v$ with $x_{b,v}=0$,
  • followed by at most one value $v$ with $0<x_{b,v}<q_{b,v}$ (the crucial value),
  • followed by a (possibly empty) suffix of values $v$ with $x_{b,v}=q_{b,v}$ (the tight values).
Per Set S.
A central challenge in our mechanism is to offer identical price menus to all buyers. While identical pricing is crucial for maintaining anonymity, it must be combined with additional structural properties, such as subadditivity across sets, which we address in later sections. In this subsection, we focus on the implications of identical pricing. Specifically, anonymity allows us to treat all buyers interested in the same set $S$ uniformly under the mechanism.
For a set $S$, let $\mathcal{A}_{S}\coloneq\{b:S_{b}=S\}$ be the set of buyers that desire $S$. For $v\in\mathcal{S}$ let $x_{S,v}=\sum_{b\in\mathcal{A}_{S}}x_{b,v}$ (define $q_{S,v}$ similarly). Denote $X^{(S)}=[x_{S,v}]_{v\in\mathcal{S}}$ (where entries are ordered in increasing order of $v$ as before). Lastly, define $\textnormal{{FracOpt}}_{S}=\sum_{v\in\mathcal{S}}x_{S,v}\cdot v$. In Appendix G we prove the following, with crucial and tight values of $S$ defined similarly to the single buyer case.
Lemma 2.4.
For each bundle $S$, the vector $X^{(S)}$ consists of three consecutive blocks:
  • a (possibly empty) prefix of values $v$ with $x_{S,v}=0$,
  • followed by at most one value $v$ with $0<x_{S,v}<q_{S,v}$ (the crucial value),
  • followed by a (possibly empty) suffix of values $v$ with $x_{S,v}=q_{S,v}$ (the tight values).
We next introduce a definition that will be essential for setting our posted prices.
Definition 2.5.
The largest value $w\in\mathcal{S}$ such that $x_{S,w}<q_{S,w}$ is called the important value of $S$ and is denoted by $w_{S}$. Note that such value always exists since we can enforce $x_{S,0}=0$. Observe also that if $S$ has a crucial value, then its important value is equal to the crucial value of $S$.
3 General Algorithmic Framework
We begin by introducing a general algorithmic framework. This framework serves as the foundation for all subsequent sections, where we instantiate it in several settings: the general single-minded case, single-minded case with maximum bundle size at most $d$ and general graphs.
3.1 A High-Level Algorithm
We present a two-step framework that converts the scaled LP solution into a static and anonymous (possibly randomized) bundle pricing. We first scale down the capacities, and then translate the per-set structure of the scaled solution into a menu of posted prices for each desired set $S$ (note that a set $S$ is desired if there exists a buyer $b$ such that $S_{b}=S$). At a high level, our framework consists of two main components.
Step 1: Scale down the LP.
We first solve the original LP but with the capacity constraints scaled down by a factor of $\gamma>1$ depending on the setting, i.e. $c(e)$ in the first LP constraint becomes $c(e)/\gamma$. Denote by $\textnormal{{FracOpt}}_{\gamma}$ the new fractional optimum. Note that $\textnormal{{FracOpt}}_{\gamma}\geq\textnormal{{FracOpt}}/\gamma$
Step 2: Post the menu M.
We now create the randomized menu $M$ from the solution of the fractional relaxation according to Algorithm˜1. Our menu is randomized in the sense that the algorithm draws a menu from a distribution before arrivals; importantly, the menu itself does not contain lotteries. The construction is guided by the per-set structure of the scaled LP. For values with $x_{S,v}=0$, we do not offer $S$ at a price that would attract these buyers. For tight values with $x_{S,v}=q_{S,v}$, we offer $|\mathcal{A}_{S}|$ copies so that we effectively accept all such buyers when they appear, with item availability (rather than menu-entry multiplicity) being the only potential bottleneck. Finally, when $S$ has a crucial value, we use a more careful (and in some cases randomized) offer at price $w$ to accept these buyers in expectation close to the mass prescribed by the LP. We are now ready to define our algorithm.
Algorithm 1 Menu Construction (Step 2)
1:Optimal solution $x$ of the scaled-down LP
2:for each (desired) bundle $S$ do
3:  Let $w$ be the important value of $S$ (w.r.t. to $x$).
4:  Offer $|\mathcal{A}_{S}|$ copies of $S$ at price $w+1$.
5:  if $S$ has crucial value $w$ and $x_{S,w}\cdot w>\displaystyle\sum_{v\in\mathcal{S}:v>w}x_{S,v}\cdot v$ then
6:   if $x_{S,w}>1$ then
7:     Offer $\lfloor x_{S,w}\rfloor$ additional copies of $S$ at price $w$.
8:   else
9:     With probability $\max\{x_{S,w},\,x_{S,w}/q_{S,w}\}$, offer one additional copy of $S$ at price $w$.
10:   end if
11:  end if
12:end for
Definition 3.1 (Algorithm).
Fix a menu $M$ and an instance of buyers’ realized values $I$. Our algorithm, denoted $\textnormal{{Alg}}_{M,I}$, operates as follows.
  1. 1.
    The adversary selects an arrival order of the buyers in $\mathcal{N}$.
  2. 2.
    Buyers arrive sequentially. When buyer $b$ arrives, they examine the current menu $M$ and choose a subset $J$ of available menu entries that maximizes their utility (possibly empty), subject to feasibility with the remaining item copies.
  3. 3.
    If $J\neq\emptyset$, then for each chosen entry $(T_{j},p_{j},c_{j})$ we decrease its remaining multiplicity by $1$, and for each item $e\in T_{j}$ we decrease the remaining copies of $e$ by $1$. Buyer $b$ is deemed accepted.
We will use $\textnormal{{Alg}}_{M,I}$ to denote either the set of buyers accepted by our algorithm or the total value they contribute; the intended meaning will be clear from the context. Our goal is to prove that $\mathbb{E}_{M,I}\left[\textnormal{{Alg}}_{M,I}\right]\geq\alpha\cdot\textnormal{{FracOpt}}$ for some $\alpha>0$.
Menu Subadditivity.
Throughout all subsequent sections, we will assume that buyers in $\mathcal{A}_{S}$ purchase only menu entries that sell exactly the bundle $S$. This is without loss of generality: in Appendix B we show that our menu prices are essentially subadditive, hence no buyer who desires $S$ can increase their utility by purchasing entries corresponding to bundles different from $S$.
Tie-Breaking.
We will also assume that buyers break ties in favour of purchasing, thus if they have value $v$ they will purchase a copy at price $v$. In Appendix B we remove this assumption with a small price perturbation.
3.2 The Unconstrained Algorithm
We now introduce an analytical tool, the Unconstrained Algorithm. This is a hypothetical version of our mechanism that completely ignores item capacities and only enforces the constraints imposed by the menu. In particular, the menu may induce total demand that exceeds the available copies of some items, and the Unconstrained Algorithm still serves all such requests. We use this algorithm purely for analysis: it is not required to be feasible, but it allows us to bound the performance of our actual (capacity-feasible) mechanism by comparing their expected welfare. More formally:
Definition 3.2 (Unconstrained Algorithm).
Fix a menu $M$ and an instance of buyers’ realized values $I$. Our unconstrained algorithm, denoted $\textnormal{{UAlg}}_{M,I}$, operates as follows.
  1. 1.
    Buyers arrive sequentially according to the order that the almighty adversary selected for $\textnormal{{Alg}}_{M,I}$. When buyer $b$ arrives, they examine the current menu $M$ and choose a subset $J$ of available menu entries that maximizes their utility (possibly empty). Item capacities are ignored.
  2. 2.
    If $J\neq\emptyset$, then for each chosen entry $(T_{j},p_{j},c_{j})$ we decrease its remaining multiplicity by $1$. Buyer $b$ is deemed accepted.
We will use $\textnormal{{UAlg}}_{M,I}$ to denote either the set of buyers accepted by the unconstrained algorithm or the total value they contribute; the intended meaning will be clear from the context. Note that a buyer that belongs to $\textnormal{{Alg}}_{M,I}$ must belong to $\textnormal{{UAlg}}_{M,I}$. We next state a key guarantee for UAlg, proved in Appendix C, showing that the unconstrained algorithm captures a constant fraction of $\textnormal{{FracOpt}}_{\gamma}$.
Lemma 3.3 (Property 1).
$\mathbb{E}_{M,I}\left[\textnormal{{UAlg}}_{M,I}\right]\geq\textnormal{{FracOpt}}_{\gamma}/8$.
Since UAlg does not take into account item capacities, it may oversell some items. Define by $L_{M,I}(e)$ the number of copies that $\textnormal{{UAlg}}_{M,I}$ sells for item $e$. In the following lemma, proved in Appendix C, we show that UAlg respects the scaled-down capacities in expectation.
Lemma 3.4 (Property 2).
For every item $e\in\mathcal{M}$ and any scaling factor $\gamma>1$, there exist $\{0,1\}$ negatively associated random variables $Z_{1},\ldots,Z_{k}$ such that
$L_{M,I}(e)\leq\sum_{i=1}^{k}Z_{i}=Z_{M,I}(e)\quad\text{and}\quad\mathbb{E}_{M,I}\left[L_{M,I}(e)\right]\leq\mathbb{E}_{M,I}\left[Z_{M,I}(e)\right]\leq\frac{c(e)}{\gamma}.$
4 d-Single-Minded Setting
In this section, we apply our framework to the case in which every buyer demands a single bundle $S_{b}\subseteq\mathcal{M}$ of size at most $d$. We proceed to prove Theorem˜1.1.
Approach.
We follow the general algorithmic framework from Section˜3. We begin by solving the scaled down LP from (Step 1), choosing $\gamma\;=\;e(10d)^{1/B}$. We then construct the randomized menu as in (Step 2).
Analysis.
Lemma 3.4 together with our choice of $\gamma=e(10d)^{1/B}$ immediately yields the following corollary.
Corollary 4.1.
For every item $e\in\mathcal{M}$,
$\mathbb{E}_{M,I}\left[Z_{M,I}(e)\right]\;\leq\;\frac{c(e)}{\gamma}\;=\;\frac{c(e)}{e(10d)^{1/B}}.$
We now show that, for every item $e\in\mathcal{M}$, the probability that its load exceeds its capacity is at most $1/(10d)$. Recall from Corollary 4.1 that
$\mathbb{E}_{M,I}\left[Z_{M,I}(e)\right]\;\leq\;\frac{c(e)}{\gamma},\qquad\text{where }\gamma=e(10d)^{1/B}.$
Since $Z_{M,I}(e)$ is a sum of negatively associated binary random variables, as shown in the proof of Lemma˜3.4, we apply the standard Chernoff bound. Writing $\mu=\mathbb{E}_{M,I}\left[Z_{M,I}(e)\right]$, for every $e\in\mathcal{M}$ we obtain
$\mathbb{P}\!\left[Z_{M,I}(e)\geq c(e)\right]\;\leq\;\left(\frac{e\mu}{c(e)}\right)^{c(e)}\;\leq\;\left(\frac{e}{\gamma}\right)^{c(e)}.$
Substituting our choice of $\gamma$ yields $\mathbb{P}\!\left[Z_{M,I}(e)\geq c(e)\right]\;\leq\;\frac{1}{10d}$ and from union bound we get for every desired set $S$
$\mathbb{P}\!\left[\exists e\in S:\,Z_{M,I}(e)\geq c(e)\right]\leq\sum_{e\in S}\mathbb{P}\!\left[Z_{M,I}(e)\geq c(e)\right]\leq|S|\cdot\frac{1}{10d}\leq 0.1,$
since $|S|\leq d$. When a buyer arrives, they may prefer some menu entry, but the algorithm may be forced to reject them because one of the items in that bundle has exhausted its capacity. In this case, we say that the buyer is rejected by the algorithm. Let
$\textsc{Blocked}_{M,I}=\{b\in\mathcal{N}:b\in\textnormal{{UAlg}}_{M,I}\text{ and }b\notin\textnormal{{Alg}}_{M,I}\},$
be the set of rejected buyers and $\textnormal{{Alg}}_{M,I}$ the set and the total value of the buyers that our algorithm gets under realization $I$ and menu $M$. We begin with the basic decomposition (note that UAlg contains Alg):
$\mathbb{E}_{M,I}\left[\textnormal{{Alg}}_{M,I}\right]\;=\;\mathbb{E}_{M,I}\left[\textnormal{{UAlg}}_{M,I}\right]\;-\;\mathbb{E}_{M,I}\left[\sum_{b\in\textsc{Blocked}_{M,I}}v_{b}\right]$ ($\star$)
We now prove the following lemma that bounds the lost value due to rejected buyers and leads to our desired approximation ratio.
Lemma 4.2.
We have
$\mathbb{E}_{M,I}\left[\sum_{b\in\textsc{Blocked}_{M,I}}v_{b}\right]\;\leq\;0.1\;\textnormal{{FracOpt}}_{\gamma}.$
Proof.
Fix a desired set $S$ with important value $w=w_{S}$ and denote by $\mathcal{A}=\mathcal{A}_{S}$ the set of buyers that desire $S$. Note that a buyer $b$ belongs to $\textsc{Blocked}_{M,I}$ if it belongs to $\textnormal{{UAlg}}_{M,I}$ and at the same time due to limited item capacities $\textnormal{{Alg}}_{M,I}$ cannot accept it. Denote by $\mathcal{E}_{b}$ the latter event where due to capacity constraints $\textnormal{{Alg}}_{M,I}$ cannot accept buyer $b$. It holds that
$\mathbb{E}_{M,I}\left[\sum_{b\in\mathcal{A}\cap\textsc{Blocked}_{M,I}}v_{b}\right]=\mathbb{E}_{M,I}\left[\sum_{b\in\mathcal{A}}v_{b}\mathbbm{1}{\left\{b\in\textsc{Blocked}_{M,I}\right\}}\right]$
[…]
Observe that any buyer with value strictly larger than $w$ is always included in UAlg. In contrast, buyers whose value equals $w$ may or may not belong to UAlg, depending on the arrival order chosen by the adversary. Thus we consider the contribution from buyers with tight and non-tight values separately.
We first focus on the second term in the last inequality. We consider three cases regarding the menu entries at price $w$ we post for set $S$. In the following we carefully handle the events $\mathbbm{1}{\left\{\mathcal{E}_{b}\right\}}$ and $\mathbbm{1}{\left\{b\in\textnormal{{UAlg}}_{M,I}\right\}}$ to decouple the randomness in the realization and the menu.
Case 1: We offer one copy with probability $x_{S,w}/q_{S,w}$.
Thus, if $\in\textnormal{{UAlg}}_{M,I}, bb$ got the single copy at crucial value and we get (recall that $\ell_{A}(e)$ denotes the load of item $e$ due to buyers in $A$):
$\mathcal{E}_{b}\subseteq\left\{\exists e\in S\,:\,\ell_{\textnormal{{UAlg}}_{M,I}\setminus\mathcal{A}}(e)+\sum_{b^{\prime}\in\mathcal{A},\ b^{\prime}\neq b}\mathbbm{1}{\left\{v_{b^{\prime}}>w\right\}}\geq c(e)\right\}\eqqcolon\mathcal{F}_{b}$
Note that (from the proof of Lemma 3.4) $\mathbb{P}\!\left[\mathcal{F}_{b}\right]\leq 0.1$ since for each $e\in S,\ M,\ I$
$\ell_{\textnormal{{UAlg}}_{M,I}\setminus\mathcal{A}}(e)+\sum_{b^{\prime}\in\mathcal{A},\ b^{\prime}\neq b}\mathbbm{1}{\left\{v_{b^{\prime}}>w\right\}}\leq Z_{M,I}(e).$
Then, the second term is upper bounded by
$\mathbb{E}_{M,I}\left[\sum_{b\in\mathcal{A}}w\mathbbm{1}{\left\{v_{b}=w\right\}}\mathbbm{1}{\left\{b\in\textnormal{{UAlg}}_{M,I}\right\}}\mathbbm{1}{\left\{\mathcal{F}_{b}\right\}}\right]$
[…]
Case 2: We offer one copy with probability $x_{S,w}$.
Define
$\mathcal{F}_{A}\coloneqq\left\{\exists e\in S\,:\,\ell_{\textnormal{{UAlg}}_{M,I}\setminus\mathcal{A}}(e)+\sum_{b\in\mathcal{A}}\mathbbm{1}{\left\{v_{b}>w\right\}}\geq c(e)\right\}$
Note that $\mathbb{P}\!\left[\mathcal{F}_{A}\right]\leq 0.1$ similarly to Case 1. Now it holds that for every $M$ and for every II
[…]
Case 3: We offer $\lfloor x_{S,w}\rfloor$ copies.
Define
$\mathcal{F}_{\mathcal{A}}\coloneqq\left\{\exists e\in S\,:\,\ell_{\textnormal{{UAlg}}_{M,I}\setminus\mathcal{A}}(e)+\sum_{b\in\mathcal{A}}\mathbbm{1}{\left\{v_{b}>w\right\}}+\left\lfloor x_{S,w}\right\rfloor\geq c(e)\right\}$
Note that $\mathbb{P}\!\left[\mathcal{F}_{\mathcal{A}}\right]\leq 0.1$ similarly to Case 1. Now it holds that for every $M$ and for every II
$\sum_{b\in\mathcal{A}}\mathbbm{1}{\left\{v_{b}=w\right\}}\mathbbm{1}{\left\{b\in\textnormal{{UAlg}}_{M,I}\right\}}\mathbbm{1}{\left\{\mathcal{E}_{b}\right\}}\leq\ \left\lfloor x_{S,w}\right\rfloor\cdot\mathbbm{1}{\left\{\mathcal{F}_{\mathcal{A}}\right\}}$
Indeed, the left-hand side is upper bounded by $\left\lfloor x_{w}\right\rfloor$ and if $\mathbbm{1}{\left\{\mathcal{F}_{\mathcal{A}}\right\}}=0$ then no buyer in $\mathcal{A}$ can be blocked. Taking expectations and multiplying with $w$ yields that the second term is upper bounded by $0.1\cdot w\left\lfloor x_{S,w}\right\rfloor\leq 0.1\cdot w\cdot x_{S,w}$.
Combining all the previous cases and substituting back to ($\star\star$4) we get:
$\mathbb{E}_{M,I}\left[\sum_{b\in\mathcal{A}\cap\textsc{Blocked}_{M,I}}v_{b}\right]\leq\mathbb{E}_{M,I}\left[\sum_{b\in\mathcal{A}}v_{b}\mathbbm{1}{\left\{v_{b}>w\right\}}\mathbbm{1}{\left\{\mathcal{E}_{b}\right\}}\right]+0.1\cdot x_{S,w}\cdot w$
Now note that
$\mathcal{E}_{b}\subseteq\left\{\exists e\in S\,:\,\ell_{\textnormal{{UAlg}}_{M,I\setminus\{b\}}}(e)\geq c(e)\right\}\eqqcolon\mathcal{G}_{b}$
since the load due to previous buyers (with respect to $b$) in any item is smaller than the load due to UAlg ignoring $b$ from the instance $I$. Now noting that $\mathbb{P}\!\left[\mathcal{G}_{b}\right]\leq 0.1$ since $\ell_{\textnormal{{UAlg}}_{M,I\setminus\{b\}}}(e)$ is upper bounded by $Z_{M,I}(e)$ and using the fact that $\mathcal{G}_{b}$ is independent to $v_{b}$ we get
$\mathbb{E}_{M,I}\left[\sum_{b\in\mathcal{A}\cap\textsc{Blocked}_{M,I}}v_{b}\right]$ $\leq\mathbb{E}_{M,I}\left[\sum_{b\in\mathcal{A}}v_{b}\mathbbm{1}{\left\{v_{b}>w\right\}}\mathbbm{1}{\left\{\mathcal{G}_{b}\right\}}\right]+0.1\cdot x_{S,w}\cdot w$
$\leq\sum_{b\in\mathcal{A}}\sum_{v>w}v\cdot q_{b,v}\cdot 0.1+0.1\cdot x_{S,w}\cdot w=0.1\textnormal{{FracOpt}}_{S}$
and summing over all possible $S$ yields the proof. ∎ As a result, ($\star$4) becomes
$\mathbb{E}_{M,I}\left[\textnormal{{Alg}}_{M,I}\right]$ $\;\geq\;\mathbb{E}_{M,I}\left[\textnormal{{UAlg}}_{M,I}\right]\;-\;0.1\textnormal{{FracOpt}}_{\gamma}\geq\frac{1}{40}\cdot\textnormal{{FracOpt}}_{\gamma}$
$\;\geq\;\frac{\textnormal{{FracOpt}}}{40e(10d)^{1/B}}$ ($\star$)
5 General Single-Minded Setting
We now turn to the most general case, where buyers are single-minded over arbitrary bundles. We proceed to prove Theorem˜1.2.
Approach.
Our algorithmic approach follows the same template as before: we solve the scaled-down LP (Step 1) with an appropriate choice of $\gamma$, construct the randomized menu (Step 2), and analyze the performance of the resulting posted-price mechanism.
In this setting, however, very large valuations can dominate the benchmark and invalidate the guarantees obtained by the LP-based menu alone. To handle this, we introduce an auxiliary mechanism that posts a single high-priced bundle; the final algorithm randomly mixes between these two mechanisms.
Analysis.
We select $\gamma\;=\;e(20m)^{1/(B+1)},$ and let $\textnormal{{FracOpt}}_{\gamma}$ be the value of the scaled LP. When a buyer arrives, they may desire a menu entry but fail to obtain it due to capacity conflicts with previously accepted buyers. As before, we define:
$\textsc{Blocked}_{M,I}=\{b\in\mathcal{N}:b\in\textnormal{{UAlg}}_{M,I}\text{ and }b\notin\textnormal{{Alg}}_{M,I}\},$
and $\textnormal{{Alg}}_{M,I}$ is the total value obtained by our (constrained) algorithm. Using the unconstrained benchmark, we have the basic identity
$\mathbb{E}_{M,I}\left[\textnormal{{Alg}}_{M,I}\right]\;=\;\mathbb{E}_{M,I}\left[\textnormal{{UAlg}}_{M,I}\right]\;-\;\mathbb{E}_{M,I}\left[\sum_{b\in\textsc{Blocked}_{M,I}}v_{b}\right].$
We now bound the loss due to rejections. Partition buyers into those with moderate value and those with very large value:
$\mathbb{E}_{M,I}\left[\sum_{b\in\textsc{Blocked}_{M,I}}v_{b}\right]$ $=\mathbb{E}_{M,I}\left[\sum_{b\in\textsc{Blocked}_{M,I}}v_{b}\,\mathbbm{1}{\left\{v_{b}<2\textnormal{{FracOpt}}_{\gamma}\right\}}\right]$
$\qquad+\mathbb{E}_{M,I}\left[\sum_{b\in\textsc{Blocked}_{M,I}}v_{b}\,\mathbbm{1}{\left\{v_{b}\geq 2\textnormal{{FracOpt}}_{\gamma}\right\}}\right]$
$\leq 2\textnormal{{FracOpt}}_{\gamma}\cdot\mathbb{E}_{M,I}\left[|\textsc{Blocked}_{M,I}|\right]$
$\qquad+\mathbb{E}_{M,I}\left[\sum_{b\in\mathcal{N}}v_{b}\,\mathbbm{1}{\left\{v_{b}\geq 2\textnormal{{FracOpt}}_{\gamma}\right\}}\right]$
Bounding the number of rejected buyers.
We begin by analyzing the first term, bounding the expected number of blocked buyers.
Lemma 5.1.
We have
$\mathbb{E}_{M,I}\left[|\textsc{Blocked}_{M,I}|\right]\;\leq\;\frac{1}{20}.$
Proof.
A buyer is rejected only if some item in their bundle is oversold. Thus
$|\textsc{Blocked}_{M,I}|\;\leq\;\sum_{e\in\mathcal{M}}\sum_{k=c(e)+1}^{\infty}\mathbbm{1}{\left\{L_{M,I}(e)=k\right\}}\,(k-c(e)).$
Taking expectations and using linearity,
$\mathbb{E}_{M,I}\left[|\textsc{Blocked}_{M,I}|\right]\;\leq\;\sum_{e\in\mathcal{M}}\sum_{k=c(e)+1}^{\infty}(k-c(e))\,\mathbb{P}\!\left[L_{M,I}(e)=k\right].$
By Lemma˜3.4, the random load $L_{M,I}(e)$ is upper bounded by a sum of negatively associated Bernoulli variables with total mean $\mu_{e}\leq c(e)/\gamma$. A standard Chernoff bound for negatively associated variables gives
$\mathbb{P}\!\left[L_{M,I}(e)\geq k\right]\;\leq\;\mathbb{P}\!\left[Z_{M,I}(e)\geq k\right]\;\leq\;\left(\frac{e\,\mu_{e}}{k}\right)^{k}\;\leq\;\left(\frac{e\,c(e)}{\gamma\,k}\right)^{k}.$
Let $k=c(e)+j$ with $j\geq 1$. Using $(1+\tfrac{j}{c})^{c+j}\geq e^{j}$ and our choice of $\gamma$, one obtains
$\left(\frac{e\,c(e)}{\gamma(c(e)+j)}\right)^{c(e)+j}\;\leq\;\left(\frac{e}{\gamma}\right)^{c(e)+j}e^{-j}\;\leq\;\frac{1}{20m}\,e^{-j}.$
Summing over $j$ gives
$\sum_{k=c(e)+1}^{\infty}(k-c(e))\,\Pr[L_{M,I}(e)=k]\;\leq\;\frac{1}{20m}.$
Summing over all $m$ items yields the bound. ∎
Plugging Lemma˜5.1 into our decomposition and using Lemma˜3.3, we obtain:
$\mathbb{E}_{M,I}\left[\textnormal{{Alg}}_{M,I}\right]\;\geq\;\frac{1}{40}\,\textnormal{{FracOpt}}_{\gamma}-\sum_{b\in\mathcal{N}}\mathbb{E}_{M,I}\left[v_{b}\,\mathbbm{1}{\left\{v_{b}\geq 2\textnormal{{FracOpt}}_{\gamma}\right\}}\right].$
The second term corresponds to buyers with extremely large valuations. We now control this contribution by mixing with a second, high-price mechanism.
Handling very large valuations.
Define an auxiliary mechanism $\textnormal{{Alg}}_{2}$ that posts a single menu entry consisting of all items, priced at $2\textnormal{{FracOpt}}_{\gamma}$. We first show that we may assume that extremely large valuations are rare.
Lemma 5.2.
$\mathbb{P}\!\left[\exists b:v_{b}\geq 2\,\textnormal{{FracOpt}}_{\gamma}\right]\leq 1/2.$
Proof.
If instead $\mathbb{P}\!\left[\exists b:v_{I}(b)\geq 2\textnormal{{FracOpt}}_{\gamma}\right]>1/2$, then simply posting one copy of the whole item set priced at $2\textnormal{{FracOpt}}_{\gamma}$ would yield expected value $\textnormal{{FracOpt}}_{\gamma}$, giving an $O(\gamma)$-approximation directly, which is already $O(m^{1/(B+1)})$. Thus we may assume the stated bound. ∎
We now show that $\textnormal{{Alg}}_{2}$ captures a constant fraction of the total contribution of very large valuations.
Lemma 5.3.
$\mathbb{E}_{M,I}\left[\textnormal{{Alg}}_{2}\right]\;\geq\;\frac{1}{2}\,\sum_{b\in\mathcal{N}}\mathbb{E}_{M,I}\left[v_{b}\,\mathbbm{1}{\left\{v_{b}\geq 2\textnormal{{FracOpt}}_{\gamma}\right\}}\right].$
Proof.
For fixed $M,I$,
[…]
Whenever all other buyers $b^{\prime}\neq b$ have $v(b^{\prime})<2\textnormal{{FracOpt}}_{\gamma}$, the copy is available for $b$. Thus
$\textnormal{{Alg}}_{2}\;\geq\;\sum_{b\in\mathcal{N}}v_{b}\,\mathbbm{1}{\left\{v_{b}\geq 2\textnormal{{FracOpt}}_{\gamma}\right\}}\,\mathbbm{1}{\left\{\forall b^{\prime}\neq b:v(b^{\prime})<2\textnormal{{FracOpt}}_{\gamma}\right\}}.$
Taking expectations and using independence of the $v_{b}$,
$\mathbb{E}_{I}\left[\textnormal{{Alg}}_{2}\right]$ $\;\geq\;\sum_{b}\mathbb{E}_{I}\left[v_{b}\mathbbm{1}{\left\{v_{b}\geq 2\textnormal{{FracOpt}}_{\gamma}\right\}}\right]\,\mathbb{P}\!\left[\forall b^{\prime}:v(b^{\prime})<2\textnormal{{FracOpt}}_{\gamma}\right]$
$\;\geq\;\frac{1}{2}\sum_{b}\mathbb{E}_{I}\left[v_{b}\mathbbm{1}{\left\{v_{b}\geq 2\textnormal{{FracOpt}}_{\gamma}\right\}}\right],$
where the last inequality applies Lemma˜5.2. ∎
Combined algorithm and guarantee.
Let $\textnormal{{Alg}}_{1}$ denote the mechanism generated by the LP-based menu, and let the final algorithm Alg run $\textnormal{{Alg}}_{1}$ with probability $1/3$ and $\textnormal{{Alg}}_{2}$ with probability $2/3$. Combining the previous bounds, we obtain
$\mathbb{E}\left[\textnormal{{Alg}}\right]$ $\geq\frac{1}{3}\left(\frac{1}{40}\textnormal{{FracOpt}}_{\gamma}-\sum_{b}\mathbb{E}\left[v_{b}\mathbbm{1}{\left\{v_{b}\geq 2\textnormal{{FracOpt}}_{\gamma}\right\}}\right]\right)$
$\qquad+\frac{2}{3}\cdot\frac{1}{2}\sum_{b}\mathbb{E}\left[v_{b}\mathbbm{1}{\left\{v_{b}\geq 2\textnormal{{FracOpt}}_{\gamma}\right\}}\right]$
$=\frac{1}{120}\textnormal{{FracOpt}}_{\gamma}\;\geq\;\frac{\textnormal{{FracOpt}}}{120e(20m)^{1/(B+1)}}$
6 Graph Setting
In this section, we extend our framework to the setting where items correspond to edges and buyers correspond to routing requests in a general graph. Let $G=(V,E)$ be an undirected graph equipped with an integer capacity function $c:E\to\mathbb{Z}_{>0}$. Each buyer $b\in[n]$ is associated with a publicly known source-destination pair $(s_{b},t_{b})$. Upon arrival, buyer $b$ draws a value $v_{b}$ from a known distribution $\mathcal{D}_{b}$. Buyer $b$ derives value $v_{b}$ if and only if they receive a bundle of edges that contains a path connecting $s_{b}$ to $t_{b}$; otherwise their utility is zero. We proceed to prove Theorem˜1.3.
Let $\mathcal{P}_{s,t}$ be the set of all simple $s$-$t$ paths in $G$ and let $\mathcal{P}_{b}=\mathcal{P}_{s_{b},t_{b}}$. For each $b$, let $q_{b,v}=\mathbb{P}\!\left[v_{b}=v\right]$. We write $p\ni e$ to denote that a path $p$ uses edge $e$. As in the single-minded setting, we compare against the optimal prophet via a fractional relaxation. The LP now contains one variable for every possible routing choice:
$\max$ $\sum_{b\in[n]}\sum_{v\in\mathcal{S}}\sum_{p\in\mathcal{P}_{b}}x_{b,v,p}\cdot v$
s.t. $\sum_{b\in[n]}\sum_{v\in\mathcal{S}}\sum_{\begin{subarray}{c}p\in\mathcal{P}_{b}:\\ e\in p\end{subarray}}x_{b,v,p}\;\leq\;c(e),$ $\forall e\in E,$
$\sum_{p\in\mathcal{P}_{b}}x_{b,v,p}\;\leq\;q_{b,v},$ $\forall b\in[n],\,v\in\mathcal{S},$
$x_{b,v,p}\geq 0,$ $\forall b\in[n],\,v\in\mathcal{S},\,p\in\mathcal{P}_{b}.$
Let FracOpt denote the value of the optimal fractional LP solution.
Buyer Types and Generalized Bundles.
A key observation is that any $s$-$t$ path satisfies all buyers with the same source-target pair $(s,t)$. This motivates treating each pair $(s,t)$ as a type whose feasible bundles are exactly the $s$-$t$ paths $\mathcal{P}_{s,t}$.
We introduce a generalized bundle $B_{s,t}$, representing the act of serving an $(s,t)$-buyer. Each copy of $B_{s,t}$ will later be instantiated as a specific simple path $p\in\mathcal{P}_{s,t}$; different copies may correspond to different paths. The path associated with each copy is sampled before any buyer arrives, from a distribution over $\mathcal{P}_{s,t}$ that we define later.
For the moment, assume that every buyer $b$ with $(s_{b},t_{b})=(s,t)$ can only purchase copies of $B_{s,t}$ [^1] In Section D.2 we remove this assumption by proving that the menu prices are subadditive, so buyers never have an incentive to purchase a collection of tickets that indirectly covers their path instead of $B_{s,t}$ itself.. Under this assumption, buyers of the same type behave exactly like single-minded buyers for a generalized bundle $B_{s,t}$. We may therefore group them as a single demand class and randomize over how many copies of $B_{s,t}$ to create, just as we randomize over bundle copies in the single-minded setting.
Crucially, all arguments that relate the LP to the expected value of the unconstrained algorithm (Property 1) depend only on the total fractional mass assigned to each “bundle” and not on its internal structure. After grouping buyers by type in this way, the analysis of UAlg from the single-minded case carries over without change. What remains is to specify how we (i) construct the menu for each type $(s,t)$ and (ii) route buyers who purchase $B_{s,t}$ so that the load bounds (Property 2) are preserved at the edge level.
Menu Construction and Routing.
We now describe the construction of the menu $M$ in the graph setting. For each pair $(s,t)$ with at least one buyer, let
$\mathcal{A}_{s,t}\coloneqq\{b\in[n]:(s_{b},t_{b})=(s,t)\}$
be the set of buyers of type $(s,t)$, and define
$x_{(s,t),v}\;=\;\sum_{b\in\mathcal{A}_{s,t}}\sum_{p\in\mathcal{P}_{s,t}}x_{b,v,p}\quad\text{and}\quad x_{(s,t),v,p}\;=\;\sum_{b\in\mathcal{A}_{s,t}}x_{b,v,p}.$
As in Section˜2.2, we can view $X^{(s,t)}=[x_{(s,t),v}]_{v\in\mathcal{S}}$ as the aggregate LP allocation over values for type $(s,t)$. Let $w_{s,t}$ be the important value of this vector, defined exactly as in the single-minded case. Depending on whether $w_{s,t}$ is tight or crucial, we apply the same three-case rule as in (Step 2) to determine how many copies of $B_{s,t}$ to post at price $w_{s,t}$ and at price $w_{s,t}+1$.
Formally, the menu for type $(s,t)$ consists of $c_{s,t}$ copies of $B_{s,t}$, each priced either at $w_{s,t}$ or $w_{s,t}+1$, according to the usual case distinction from Section˜3.1. This fully specifies the prices and multiplicities of the generalized bundles.
We now define how each copy of $B_{s,t}$ is instantiated as a concrete $s$-$t$ path. For every copy of $B_{s,t}$ priced at $w$, we independently sample a path $p\in\mathcal{P}_{s,t}$ using roulette drawing with probabilities proportional to the LP mass at value $w$ (we omit the $(s,t)$ index, since we fix an $(s,t)$ pair):
$p\sim\frac{x_{w,p}}{\sum_{p^{\prime}\in\mathcal{P}_{s,t}}x_{w,p^{\prime}}}.$
Similarly, for every copy of $B_{s,t}$ priced at $w+1$, we independently sample
$p\sim\frac{\sum_{v>w}x_{v,\,p}}{\sum_{p^{\prime}\in\mathcal{P}_{s,t}}\sum_{v>w}x_{v,\,p^{\prime}}}.$
All copies are sampled independently of each other. Once the menu is fixed, when a buyer of type $(s,t)$ purchases a copy of $B_{s,t}$, they receive the pre-sampled path associated with that copy.
This completes the definition of the menu and routing policy in the graph setting.
Load Analysis.
We now show that the expected load induced by this construction satisfies the same kind of bound as in Lemma˜3.4, yielding Property 2 for graphs.
For a fixed type $(s,t)$ and edge $e$, let $L_{s,t}(e)$ denote the total load on edge $e$ due to buyers of type $(s,t)$ under the (unconstrained) execution of the menu. Let $P_{s,t}^{e}\subseteq\mathcal{P}_{s,t}$ be the set of $s$-$t$ paths that use edge $e$. The following lemma is proved using a similar logic as in the single-minded case. A formal proof of the lemma is deferred to the Section˜D.1.
Lemma 6.1 (Property 2).
For every edge $e\in E$ and any scaling factor $\gamma>1$, there exist $\{0,1\}$ negatively associated random variables $Z_{1},\ldots,Z_{k}$ such that
$L_{M,I}(e)\leq\sum_{i=1}^{k}Z_{i}=Z_{M,I}(e)\quad\text{and}\quad\mathbb{E}_{M,I}\left[L_{M,I}(e)\right]\leq\mathbb{E}_{M,I}\left[Z_{M,I}(e)\right]\leq\frac{c(e)}{\gamma}.$
Combining Lemma˜6.1 with the value guarantee for the unconstrained algorithm (Lemma˜3.3), and with the same blocking analysis used in the single-minded setting, we obtain the claimed $O\bigl(m^{1/(B+1)}\bigr)$ competitive ratio for the graph-routing model.
7 Lower Bounds
We now formally prove our lower bounds for the general capacity setting, i.e. Theorem˜1.4 and Theorem˜1.5. The refined lower bounds for the unit capacity setting, i.e. Theorem˜1.6 and Theorem˜1.7 are deferred to Appendix˜E.
7.1 Proof of Theorem 1.4
It suffices to prove the theorem under the assumption $1\leq B<\ln m$, since for $B\geq\ln m$ the trivial $\Omega(1)$ lower bound already yields the claim. We recall the notion of $r$-way qualitatively independent (QI) partitions due to Rényi [renyi].
Definition 7.1 ($r$-way qualitatively independent partitions).
Let $U$ be an $n$-element ground set. A partition of $U$ into $t$ classes is a collection $\mathcal{P}=\{P_{1},\dots,P_{t}\}$ of disjoint sets whose union is $U$. A family of partitions
$\mathcal{F}=\{\mathcal{P}^{(1)},\dots,\mathcal{P}^{(k)}\},\qquad\mathcal{P}^{(j)}=\{P^{(j)}_{1},\dots,P^{(j)}_{t}\}$,
is called $r$-way qualitatively independent if for every choice of $r$ distinct indices $j_{1},\dots,j_{r}\in[k]$ and every choice of classes $a_{1},\dots,a_{r}\in[t]$ we have
$P^{(j_{1})}_{a_{1}}\cap\dots\cap P^{(j_{r})}_{a_{r}}\neq\emptyset$.
Let $g(n,t,r)$ denote the largest integer $k$ for which such a family of $k$ partitions of an $n$-element set into $t$ classes exists. Poljak and Tuza [poljak] proved the following lower bound.
Theorem 7.2 (Theorem 5 in [poljak]).
For all integers $n,t,r\geq 1$,
$g(n,t,r)\;\geq\;\frac{er}{t}\cdot\exp\!\left(\frac{n}{rt^{r}}\right).$
We will apply this bound with $n$ equal to the number of items $m$ and $r=B+1$. The key step is to choose the number of classes $t$ so that $g(m,t,r)$ is approximately $Bt^{t}$. In Appendix G we prove the following.
Lemma 7.3.
For every integer $m\geq 16$ and every integer $B$ with $1\leq B<\ln m$, define
$t=t_{m,B}\coloneqq\left\lfloor\left(\frac{m}{2B\ln m}\right)^{1/(B+2)}\right\rfloor.$
Then $g(m,t,B+1)\;\geq\;B\,t^{\,t}.$
Lower-Bound Instance.
We now construct the lower-bound instance in the prophet-inequality setting. Let $m\geq 16$ denote the number of items and suppose that we have exactly $B$ copies of each item. Set
$t=t_{m,B}\coloneqq\left\lfloor\left(\frac{m}{2B\ln m}\right)^{1/(B+2)}\right\rfloor,$
and let $N=Bt^{t}\leq g(m,t,B+1)$. By Lemma 7.3, there exists a $(B+1)$-way qualitatively independent family
$\mathcal{F}=\left\{\mathcal{P}^{(1)},\dots,\mathcal{P}^{(N)}\right\}$
of $N$ partitions of the $m$ items into $t$ classes each:
$\mathcal{P}^{(i)}=\bigl\{P^{(i)}_{1},\dots,P^{(i)}_{t}\bigr\},\qquad i\in[N].$
For each partition $i\in[N]$ and each class $a\in[t]$, introduce a single-minded buyer $B_{i,a}$ who wants exactly the bundle $P^{(i)}_{a}$ and whose value is
$v_{i,a}=\begin{cases}1&\text{with probability }1/t,\\ 0&\text{with probability }1-1/t,\end{cases}$
independently over all pairs $(i,a)$.
Analysis.
For a fixed $i$, the number of buyers in the $i$-th group $\{B_{i,1},\dots,B_{i,t}\}$ whose value is $1$ is distributed as $\textnormal{Bin}(t,1/t)$. For each group $i\in[N]$, let $E_{i}$ be the event that all $t$ buyers in group $i$ have value $1$. These events are independent and satisfy
$\mathbb{P}\!\left[E_{i}\right]=p=(1/t)^{t}.$
Let $X$ be the number of groups in which all $t$ buyers have value $1$; then
$X\sim\textnormal{Bin}\bigl(N,(1/t)^{t}\bigr),\qquad\mathbb{E}\left[X\right]=Np=B.$
By Lemma G.3, we get $\mathbb{P}\!\left[X\geq B\right]\geq 0.5$. We next relate the combinatorial structure of the partitions to the capacity constraints via to simple observations.
Observation 7.4.
In any feasible allocation (respecting the $B$-copy constraint for each item), it is impossible to serve buyers from more than $B$ distinct groups.
Observation 7.5.
Suppose $E_{i_{1}},\dots,E_{i_{B}}$ all occur for $B$ distinct groups $i_{1},\dots,i_{B}$. Then there exists a feasible allocation that serves all $tB$ buyers in these $B$ groups.
We now bound the offline and online performance.
Offline Prophet.
Whenever $X\geq B$, by Observation 7.5 the offline optimum can choose any $B$ groups $i_{1},\dots,i_{B}$ among those with $E_{i_{j}}$ and serve all $t$ buyers in each of them, obtaining value $Bt$. Thus
$\mathrm{OPT}\;\geq\;Bt\cdot\mathbbm{1}{\left\{\{X\geq B\}\right\}},$
and hence
$\mathbb{E}\left[\mathrm{OPT}\right]\;\geq\;Bt\cdot\mathbb{P}\!\left[X\geq B\right]\;\geq\;0.5\,Bt$
for all sufficiently large $m$.
Online Algorithm.
Consider an arbitrary online algorithm. For each group $i\in[N]$, let $Y_{i}$ be the number of buyers in group $i$ with value $1$. Then $Y_{i}\sim\textnormal{Bin}(t,1/t)$ and thus $\mathbb{E}\left[Y_{i}\right]=1$. By Observation 7.4, the algorithm can serve buyers from at most $B$ distinct groups. Let $t_{i}$ denote the $i$-th different group (chronologically) that the algorithm selects to serve, thus:
$\mathbb{E}\left[\mathrm{ALG}\right]\;\leq\;\sum_{i=1}^{B}\mathbb{E}\left[\text{value from group $t_{i}$}\right]\;\leq\;2B.$
since after we select a buyer from group $t_{i}$ the expected total value of the remaining buyers from this group is less than $1$.
Competitive ratio.
Combining the two bounds, for all sufficiently large $m$,
$\frac{\mathbb{E}[\mathrm{OPT}]}{\mathbb{E}[\mathrm{ALG}]}\;\geq\;\frac{0.5\,Bt}{2B}=0.25\,t.$
Recall that
$t=\left\lfloor\left(\frac{m}{2B\ln m}\right)^{1/(B+2)}\right\rfloor$
For $m\geq 16$, the term in the floor is at least $1$, so
$t\;\geq\;\frac{1}{2}\left(\frac{m}{2B\ln m}\right)^{1/(B+2)}\;=\;\frac{1}{2}\,\left(\frac{1}{2B}\right)^{1/(B+2)}\left(\frac{m}{\ln m}\right)^{1/(B+2)}\geq\frac{1}{2}\,\cdot\frac{1}{2}\cdot\left(\frac{m}{\ln m}\right)^{1/(B+2)}.$
Thus:
$\frac{\mathbb{E}[\mathrm{OPT}]}{\mathbb{E}[\mathrm{ALG}]}\;\geq\;\frac{1}{16}\left(\frac{m}{\ln m}\right)^{1/(B+2)}.$
which proves Theorem˜1.4.
7.2 Proof of Theorem 1.5
We give a variant of the lower-bound template from Theorem˜1.4 that additionally enforces a maximum desired bundle size $d$. First we introduce the notion of balanced $r$-way QI partitions.
Balanced r-way QI Partitions.
Let $U$ be an $n$-element ground set. A partition of $U$ into $t$ classes is called $d$-balanced if every class has size at most $d$. Let $g_{\mathrm{bal}}(n,t,r,d)$ denote the largest size of an $r$-way QI family of $d$-balanced $t$-partitions of an $n$-element set.
We now refine the probabilistic argument of [poljak] to the balanced setting. The only difference is that we sample partitions from uniformly random balanced labelings.
Lemma 7.6 (Balanced refinement of Theorem 4 in [poljak]).
Let $n,t,r\geq 1$ be integers with $t\mid n$, and set $d:=n/t$. Then
$g_{\mathrm{bal}}(n,t,r,d)\;\geq\;\frac{e\,r}{t}\cdot\exp\!\left(\frac{n}{rt^{r}}\right),$
Proof.
Let $U$ be an $n$-element set with $n=dt$. We generate a random $d$-balanced $t$-partition by choosing a uniformly random labeling $\sigma:U\to[t]$ such that each label appears exactly $d$ times, and letting the classes be $\sigma^{-1}(1),\dots,\sigma^{-1}(t)$. Sample $k$ such partitions independently.
Fix distinct indices $j_{1},\dots,j_{r}\in[k]$ and classes $a_{1},\dots,a_{r}\in[t]$. Let
$Z:=\Bigl|P^{(j_{1})}_{a_{1}}\cap\cdots\cap P^{(j_{r})}_{a_{r}}\Bigr|=\sum_{u\in U}\mathbbm{1}{\left\{u\in P^{(j_{1})}_{a_{1}}\cap\cdots\cap P^{(j_{r})}_{a_{r}}\right\}}$.
For each $u\in U$ and each $\ell\in[r]$,
$\mathbb{P}\!\left[u\in P^{(j_{\ell})}_{a_{\ell}}\right]=\frac{d}{n}=\frac{1}{t},$
and by independence across the $r$ sampled partitions,
$\mathbb{P}\!\left[u\in P^{(j_{1})}_{a_{1}}\cap\cdots\cap P^{(j_{r})}_{a_{r}}\right]=\frac{1}{t^{r}}.$
Hence $\mathbb{E}\left[Z\right]=\frac{n}{t^{r}}$. Moreover, for a fixed partition $j$ and class $a$, $P^{(j)}_{a}$ is a uniformly random subset of $U$ of size $d$. It is a well known fact that for a uniformly random size-$d$ subset, the membership indicators are negatively associated (sampling without replacement) therefore
$\mathbb{P}\!\left[Z=0\right]\;\leq\;\prod_{u\in U}\mathbb{P}\!\left[u\notin P^{(j_{1})}_{a_{1}}\cap\cdots\cap P^{(j_{r})}_{a_{r}}\right]=\left(1-\frac{1}{t^{r}}\right)^{n}$
Let $\mathrm{Bad}$ be the event that the $k$ sampled partitions are not $r$-way QI. By a union bound over the $\binom{k}{r}$ choices of distinct indices and the $t^{r}$ choices of classes,
$\mathbb{P}\!\left[\mathrm{Bad}\right]\;\leq\;\binom{k}{r}\,t^{r}\left(1-\frac{1}{t^{r}}\right)^{n}.$
Similarly to [poljak], choosing
$k=\left\lceil\frac{er}{t}\cdot\exp\!\left(\frac{n}{rt^{r}}\right)\right\rceil$
ensures $\mathbb{P}\!\left[\mathrm{Bad}\right]<1$, hence proving the lemma. ∎
We now set $r:=B+1$, since $(B+1)$-way QI is exactly what rules out serving buyers from more than $B$ groups when there are $B$ copies of each item (as in Observation 7.4). We will take $m=dt$ items and require that every class in our partitions has size at most $d$; with $t\mid m$, balanced sampling gives classes of size exactly $m/t=d$.
Define, for $d$ sufficiently large and $1\leq B<\ln d$,
$t=t_{d,B}\;\coloneqq\;\left\lfloor\left(\frac{d}{2B\ln d}\right)^{\!\frac{1}{B+1}}\right\rfloor,\qquad m:=dt.$ (1)
In Appendix G we prove the following.
Lemma 7.7.
With $t$ and $m$ as in (1),
$g_{\mathrm{bal}}(m,t,B+1,d)\;\geq\;B\,t^{\,t},$
for all sufficiently large $d$.
Lower-Bound Instance with Bundle Cap d.
Let $m=dt$ items and suppose there are exactly $B$ copies of each item. By Lemma˜7.7, there exists a $(B+1)$-way QI family
$\mathcal{F}=\bigl\{\mathcal{P}^{(1)},\dots,\mathcal{P}^{(N)}\bigr\},\qquad N:=Bt^{t},$
of $d$-balanced partitions of the $m$ items into $t$ classes each:
$\mathcal{P}^{(i)}=\{P^{(i)}_{1},\dots,P^{(i)}_{t}\},\qquad|P^{(i)}_{a}|\leq d.$
For each $i\in[N]$ and $a\in[t]$, introduce a single-minded buyer $B_{i,a}$ who wants exactly $P^{(i)}_{a}$ and whose value is
$v_{i,a}=\begin{cases}1&\text{with probability }1/t,\\ 0&\text{with probability }1-1/t,\end{cases}$
independently over all $(i,a)$. Thus every desired bundle has size at most $d$.
Analysis.
Define events $E_{i}$ and the random variable $X$ exactly as in Theorem˜1.4. Then $X\sim\mathrm{Bin}(N,(1/t)^{t})$ and $Np=B$, so
$\mathbb{P}\!\left[X\geq B\right]\geq 0.5$
by Lemma˜G.3.
The structural claims from Theorem˜1.4 remain unchanged:
  • By $(B+1)$-way QI, no feasible allocation can serve buyers from more than $B$ distinct groups.
  • If $E_{i_{1}},\dots,E_{i_{B}}$ occur, all $tB$ buyers in these $B$ groups can be served.
Therefore,
$\mathbb{E}\left[\mathrm{OPT}\right]\;\geq\;0.5\,Bt,\qquad\mathbb{E}\left[\mathrm{ALG}\right]\;\leq\;2B,$
and hence
$\frac{\mathbb{E}\left[\mathrm{OPT}\right]}{\mathbb{E}\left[\mathrm{ALG}\right]}\;\geq\;0.25\,t.$
Competitive Ratio as a function of d and B.
From (1), for all sufficiently large $d$,
$t\;\geq\;\frac{1}{2}\left(\frac{d}{2B\ln d}\right)^{\!\frac{1}{B+1}},$
and thus every online algorithm has competitive ratio at least
$\Omega\!\left(\left(\frac{d}{\ln d}\right)^{\!\frac{1}{B+1}}\right).$
8 Conclusion
We introduced a unified approach for online resource allocation with strong complementarities via static and anonymous bundle-pricing. Our mechanisms are guided by an ex-ante LP relaxation with scaled down capacities, and translate the fractional solution’s threshold structure into posted bundle prices. This yields improved guarantees as multiplicity grows: an $O(d^{1/B})$-competitive mechanism for the $d$-single-minded setting, and an $O(m^{1/(B+1)})$-competitive mechanism for general single-minded combinatorial auctions and for capacitated graph routing. We also prove information-theoretic lower bounds via an embedding of the classical prophet-inequality hard instance using qualitatively-independent partitions, showing that the dependence on $m$ (or $d$) and $B$ is essentially unavoidable. Two directions seem particularly interesting. First, closing the remaining gaps for both the general and $d$-single-minded setting. Second, sample-based implementations: our mechanism assumes that the value distributions of the buyers are explicitly given to solve the ex-ante LP and set prices. Can one obtain comparable guarantees using few samples?
Appendix A Different Distribution Supports
Our results extend to the setting where each buyer $b$ has a distribution $\mathcal{D}_{b}$ over its own support $\mathcal{S}_{b}$ and by a standard scaling argument we can assume that $\mathcal{S}_{b}\subseteq\mathbb{Z}_{\geq 0}$ without loss of generality. We extend each buyer-specific support $\mathcal{S}_{b}$ to a common support $\mathcal{S}$, run our $c$-approximation algorithm in the resulting same-support instance, and post the induced randomized menu. We show that with the appropriate extension, this (randomized) menu yields a $4c/3$-approximation for the original problem with different supports.
Let $v_{\max}^{(b)}=\max_{v\in\mathcal{S}_{b}}v$ and $v_{\max}=\max_{b}v_{\max}^{(b)}$. Define $\mathcal{S}=\{0,\ldots,v_{\max}\}\supseteq\cup_{b}\mathcal{S}_{b}$. To reduce to the equal-support case, we modify each distribution $\mathcal{D}_{b}$ so that all of them share the same support $\mathcal{S}$. For each buyer $b$ with $\mathcal{S}\setminus\mathcal{S}_{b}\neq\emptyset$, define the modified distribution $\mathcal{D}_{b}^{\prime}$ by
$q^{\prime}_{b,v^{(b)}_{\max}}\;=\;q_{b,v^{(b)}_{\max}}-\varepsilon\cdot|\mathcal{S}\setminus\mathcal{S}_{b}|,\qquad q^{\prime}_{b,v}=\varepsilon\quad\forall v\in\mathcal{S}\setminus\mathcal{S}_{b},$
and leave all other $q_{b,v}$ unchanged. The parameter $\varepsilon>0$ is chosen sufficiently small so that
$0<q^{\prime}_{b,v^{(b)}_{\max}}\quad\text{i.e. }\quad\varepsilon\leq C_{1}\coloneqq\min_{b}\frac{q_{b,v^{(b)}_{\max}}}{2\,|\mathcal{S}\setminus\mathcal{S}_{b}|}.$
The modified distribution $\mathcal{D}_{b}^{\prime}$ is a valid distribution, and all $\mathcal{D}_{b}^{\prime}$ share common support $\mathcal{S}$. Define $\mathcal{D}\coloneqq\mathcal{D}_{1}\times\ldots\times\mathcal{D}_{n}$ and $\mathcal{D^{\prime}}\coloneqq\mathcal{D}^{\prime}_{1}\times\ldots\times\mathcal{D}^{\prime}_{n}$.
Let FracOpt and $\textnormal{{FracOpt}}^{\prime}$ be the fractional LP optima under the original and modified distributions, respectively. Now take an optimal LP solution $x$ for $\mathcal{D}$ and decrease $x_{b,v^{(b)}_{\max}}$ by
$\min\left\{x_{b,v^{(b)}_{\max}},\varepsilon|\mathcal{S}\setminus\mathcal{S}_{b}|\right\}$
for each buyer $b$; the resulting solution is feasible for $\mathcal{D}^{\prime}$ and loses at most $\varepsilon|\mathcal{S}\setminus\mathcal{S}_{b}|\,v_{\max}^{(b)}$ in value for each buyer $b$. Thus,
$\textnormal{{FracOpt}}^{\prime}\geq\textnormal{{FracOpt}}\;-\;\varepsilon\sum_{b}|\mathcal{S}\setminus\mathcal{S}_{b}|\,v^{(b)}_{\max}$.
Thus, by choosing
$\varepsilon\leq C_{2}\coloneqq\min\!\left\{C_{1},\;\frac{\textnormal{{FracOpt}}}{2\sum_{b}|\mathcal{S}\setminus\mathcal{S}_{b}|\,v^{(b)}_{\max}}\right\}$,
we guarantee $\textnormal{{FracOpt}}^{\prime}\geq\frac{1}{2}\textnormal{{FracOpt}}$.
Let $I\sim\mathcal{D}$ and $I^{\prime}\sim\mathcal{D}^{\prime}$ denote value realizations under the original and modified distributions. For the total variation distance between $\mathcal{D}_{b},\mathcal{D}_{b}^{\prime}$ for each buyer $b$ we get by definition that $d_{\mathrm{TV}}(\mathcal{D}_{b},\mathcal{D}_{b}^{\prime})\;=\;\varepsilon\,|\mathcal{S}\setminus\mathcal{S}_{b}|$. Thus, since $\mathcal{D},\mathcal{D}^{\prime}$ are product distributions we get
$d_{\mathrm{TV}}(\mathcal{D},\mathcal{D}^{\prime})\;\leq\;\sum_{b}d_{\mathrm{TV}}(\mathcal{D}_{b},\mathcal{D}_{b}^{\prime})\;=\;\varepsilon\sum_{b}|\mathcal{S}\setminus\mathcal{S}_{b}|.$
For any fixed menu $M$ and any realization $I$, the welfare of our algorithm is bounded as
$0\;\leq\;\textnormal{{Alg}}_{M,I}\;\leq\;\sum_{b\in\mathcal{N}}v_{\max}^{(b)}\;\eqqcolon\;V_{\max}^{\mathrm{tot}}.$
Therefore, for any fixed $M$, using a standard property of the total variation distance (see Proposition 4.5 in [levin_peres_wilmer_markovmixing]), we get
$\Bigl|\mathbb{E}_{I}\left[\textnormal{{Alg}}_{M,I}\right]-\mathbb{E}_{I^{\prime}}\left[\textnormal{{Alg}}_{M,I^{\prime}}\right]\Bigr|\;\leq\;d_{\mathrm{TV}}(\mathcal{D},\mathcal{D}^{\prime})\cdot V_{\max}^{\mathrm{tot}}\;\leq\;\varepsilon\!\left(\sum_{b}|\mathcal{S}\setminus\mathcal{S}_{b}|\right)\!V_{\max}^{\mathrm{tot}}.$
Taking expectation over the random menu $M$ yields
$\Bigl|\mathbb{E}_{M,I}\left[\textnormal{{Alg}}_{M,I}\right]-\mathbb{E}_{M,I^{\prime}}\left[\textnormal{{Alg}}_{M,I^{\prime}}\right]\Bigr|\;\leq\;\varepsilon\!\left(\sum_{b}|\mathcal{S}\setminus\mathcal{S}_{b}|\right)\!V_{\max}^{\mathrm{tot}}.$
Thus, by additionally choosing
$\varepsilon\;\leq\;C_{3}\coloneqq\min\left\{C_{2},\frac{\textnormal{{FracOpt}}}{4c\,\left(\sum_{b}|\mathcal{S}\setminus\mathcal{S}_{b}|\right)\!V_{\max}^{\mathrm{tot}}}\right\},$
we get
$\mathbb{E}_{M,I}\left[\textnormal{{Alg}}_{M,I}\right]\;\geq\;\mathbb{E}_{M,I^{\prime}}\left[\textnormal{{Alg}}_{M,I^{\prime}}\right]\;-\;\frac{\textnormal{{FracOpt}}}{4c}\;\geq\;\,\frac{\textnormal{{FracOpt}}^{\prime}}{c}\;-\;\frac{\textnormal{{FracOpt}}}{4c}.$
Combining with $\textnormal{{FracOpt}}^{\prime}\geq\frac{1}{2}\textnormal{{FracOpt}}$ gives
$\mathbb{E}_{M,I}\left[\textnormal{{Alg}}_{M,I}\right]\;\geq\;\frac{3}{4c}\,\textnormal{{FracOpt}}.$
Hence, a $c$-approximation algorithm for the equal-support case yields a $4c/3$-approximation algorithm for the general case, thus we can simply run our algorithm on the modified distributions $\mathcal{D}_{b}^{\prime}$ and post the resulting randomized menu, justifying our assumption.
Appendix B Price Subadditivity
We now relax the simplifying assumption that a buyer who desires a bundle $S$ is only interested in $S$ copies. In principle, a buyer could purchase several bundles $T_{1},\dots,T_{\ell}$ from the posted menu such that $\bigcup_{i}T_{i}\supseteq S$. We show that after a simple normalization step applied to the fractional LP solution, no buyer can ever strictly prefer such a cover over purchasing $S$ directly. In other words, the posted prices are subadditive.
Pre-processing.
After solving the scaled-down LP of Step 1, let $\mathcal{H}$ denote the set of feasible fractional solutions $x$ attaining value $\textnormal{{FracOpt}}_{\gamma}$. Among these, we select a canonical solution by solving the secondary LP
$\min\;\Bigl(\textsc{FracWeight}(x)+\sum_{e\in\mathcal{M}}w_{e}(x)\Bigr)\qquad\text{s.t. }x\in\mathcal{H},$
where
$\textsc{FracWeight}(x)=\sum_{b}\sum_{v\in\mathcal{S}}x_{b,v},\qquad w_{e}(x)=\sum_{b:e\in S_{b}}\sum_{v\in\mathcal{S}}x_{b,v}.$
We henceforth work only with optimal solutions of this normalization LP. We first prove the following lemma.
Lemma B.1.
Fix a desired set $S$ and a value $v$ such that $x_{S,v}<q_{S,v}$. Suppose there exist different desired sets $T_{1},\dots,T_{\ell}$ such that $\bigcup_{i=1}^{\ell}T_{i}\supseteq S$. If $x_{T_{i},v_{i}}>0$ for all $i\in[\ell]$ and some $v_{i}$, then
$\sum_{i=1}^{\ell}v_{i}>v.$
Proof.
Suppose, for the sake of contradiction, that $\sum_{i=1}^{\ell}v_{i}\leq v$. First, if $v=0$ the inequality trivially holds, thus we can assume that $v>0$. Then, if $\sum_{i=1}^{\ell}v_{i}<v$, since $v_{1},\dots,v_{\ell}$ are crucial or tight values and $v$ is non-tight we can transfer $\epsilon>0$ mass (for sufficiently small $\epsilon$) from each $x_{b_{i},v_{i}}>0$ to $x_{b_{j},v}$ for some choice of buyers $b_{i}\in\mathcal{A}_{T_{i}},\,b_{j}\in\mathcal{A}_{S}$ for $i\in[\ell]$, and remain feasible while increasing the objective; contradiction.
Thus suppose that $\sum_{i=1}^{\ell}v_{i}=v$. We will get contradiction due to the objective of the pre-processing step. Indeed we can do the same transfer as before and get,
  • If $\ell=1$ then $T_{1}\supsetneq S$, so the load $w_{e}(x)$ decreases by $\epsilon$ on every item $e\in T_{1}\setminus S$, while $\textsc{FracWeight}(x)$ is preserved. Thus the total objective decreases, contradiction.
  • If $\ell>1$ then the value shifted from $\ell$ different buyers is consolidated into a single buyer, so $\textsc{FracWeight}(x)$ decreases by $(\ell-1)\epsilon$, while each $w_{e}(x)$ remains the same or decreases. Again, the total objective decreases, contradiction.
Subadditivity.
Fix a buyer whose desired set is $S$. Suppose that in the posted menu they purchase bundles $T_{1},\dots,T_{\ell}$ with and
$\bigcup_{i}T_{i}\supseteq S\quad\text{and}\quad\sum_{i=1}^{\ell}p(T_{i})<p(S)$
Note that since we always offer $|\mathcal{A}_{S}|$ copies of set $S$ at some price, a copy of $S$ will always be available. Indeed, for the buyer that arrives first this trivially holds and since we will prove that the menu is subadditive they will not purchase bundles of any other desired set, thus the $|\mathcal{A}_{S}|$ are enough for each desired set $S$ (they are dedicated to buyers that desire $S$).
First, if $p(S)=1$ then $p(T_{i})=0$, contradiction. Now notice that $x_{T_{i},p(T_{i})}>0$ and $x_{S,p(S)-1}$ is not tight by the definition of our pricing rules. Thus, we can apply Lemma˜B.1 with $v_{i}=p(T_{i})$ and v=p(S)1v=p(S)-1 and get
$\sum_{i=1}^{\ell}p(T_{i})>p(S)-1,$
that contradicts with (B).
Tie-Breaking.
We now show that no tie-breaking is needed by a small price perturbation. More specifically, instead of posting price $p$ for a copy we post price $f(p)=(1-\delta)\cdot p$, for $\delta<1/(2R)$, where $R=v_{\max}$. Indeed p1<f(p)<pp-1<f(p)<p thus $p$ value buyers now gain positive utility by purchasing the copy, so no tie-breaking is needed, while buyers with value <p<p still do not afford buying the copy. Note that subadditivity of the menu still holds since the $(1-\delta)$ factor cancels out.
Appendix C Analysis of the General Algorithmic Framework
C.1 Proof of Lemma 3.3
Proof.
Fix a desired bundle $S$ and write $\mathcal{A}=\mathcal{A}_{S}$ and $w=w_{S}$ for its important value. Recall that $\textnormal{{FracOpt}}_{S}\;=\;\sum_{b\in\mathcal{A}}\sum_{v\in\mathcal{S}}x_{b,v}\,v$ is the contribution of buyers in $\mathcal{A}$ to the scaled-down fractional optimum $\textnormal{{FracOpt}}_{\gamma}$, and $x_{S,v}\;=\;\sum_{b\in\mathcal{A}}x_{b,v}$, $q_{S,v}\;=\;\sum_{b\in\mathcal{A}}q_{b,v}$. By Lemma 2.4, the vector $X^{(S)}=(x_{S,v})_{v\in\mathcal{S}}$ consists of a prefix of zeros, possibly a single crucial value $w$, and then a suffix of tight entries $x_{S,v}=q_{S,v}$ for all $v>w$. We now analyze the expected value obtained by the unconstrained algorithm UAlg from buyers in $\mathcal{A}$ under the menu of Step 2.
Case 1: S has no crucial value. Then $x_{S,v}=0$ for $v\leq w$ and $x_{S,v}=q_{S,v}$ for all $v>w$. The menu offers $|\mathcal{A}|$ copies of $S$ at price $w+1$. Since there are at most $|\mathcal{A}|$ buyers wanting $S$, no buyer is ever blocked in UAlg, and buyers with $v\leq w$ never buy at price $w+1$. Thus
$\mathbb{E}\left[\textnormal{{UAlg}}\text{ value from }S\right]\;=\;\sum_{b\in\mathcal{A}}\sum_{v>w}q_{b,v}\,v\;=\;\sum_{b\in\mathcal{A}}\sum_{v>w}x_{b,v}\,v\;=\;\textnormal{{FracOpt}}_{S}.$
Case 2: S has a crucial value w and $x_{S,w}\,w\leq\sum_{v>w}x_{S,v}\,v$. The menu is again $|\mathcal{A}|$ copies of $S$ at price $w+1$, so the same argument as in Case 1 gives
$\mathbb{E}\left[\textnormal{{UAlg}}\text{ value from }S\right]\;=\;\sum_{v>w}x_{S,v}\,v=Q_{S}.$
By assumption we get that $Q_{S}\geq\textnormal{{FracOpt}}_{S}/2$, hence
$\mathbb{E}\left[\textnormal{{UAlg}}\text{ value from }S\right]\;\geq\;\frac{1}{2}\,\textnormal{{FracOpt}}_{S}.$
Case 3: S has a crucial value w and $x_{S,w}\,w>\sum_{v>w}x_{S,v}\,v$. Now the mass at $w$ dominates. Let $Y$ denote the number of buyers in $\mathcal{A}$ whose value equals $w$; then $Y$ is a sum of independent binary random variables with total mean $q_{S,w}$. We consider two subcases, depending on $x_{S,w}$.
Case 3a: $x_{S,w}\leq 1$. The menu offers: $|\mathcal{A}|$ copies of $S$ at price $w+1$, and one additional copy at price $w$ with probability $p\;\coloneqq\;\max\{x_{S,w},\,x_{S,w}/q_{S,w}\}$. We distinguish according to $q_{S,w}$:
  • If $q_{S,w}\leq 1$, then $x_{S,w}/q_{S,w}\geq x_{S,w}$, so $p=x_{S,w}/q_{S,w}$. By Lemma G.3, $\Pr[Y\geq 1]\geq q_{S,w}/2$. Conditioned on including the $w$-priced copy and on $Y\geq 1$, UAlg sells that copy at value $w$, hence the expected contribution of this copy is at least
    $p\cdot\Pr[Y\geq 1]\cdot w\;\geq\;\frac{x_{S,w}}{q_{S,w}}\cdot\frac{q_{S,w}}{2}\cdot w\;=\;\frac{1}{2}\,x_{S,w}\,w.$
  • If $q_{S,w}>1$, then $x_{S,w}/q_{S,w}<x_{S,w}$, so $p=x_{S,w}$. Lemma G.3 yields $\Pr[Y\geq 1]\geq 1/2$, and thus the expected contribution of the $w$-priced copy is at least
    $p\cdot\Pr[Y\geq 1]\cdot w\;\geq\;x_{S,w}\cdot\frac{1}{2}\cdot w\;=\;\frac{1}{2}\,x_{S,w}\,w.$
In either subcase,
$\mathbb{E}\left[\textnormal{{UAlg}}\text{ value from }S\right]\;\geq\;\frac{1}{2}\,x_{S,w}\,w\;>\;\frac{1}{4}\,\textnormal{{FracOpt}}_{S}.$
where in the last inequality we use $x_{S,w}\,w>\sum_{v>w}x_{S,v}\,v$.
Case 3b: $x_{S,w}>1$. The menu offers: $\lfloor x_{S,w}\rfloor$ copies of $S$ at price $w$, and $|\mathcal{A}|$ copies at price $w+1$. Lemma G.3 with $\mu=q_{S,w}$ gives
$\Pr\bigl[Y\geq q_{S,w}/2\bigr]\;\geq\;\frac{1}{2},$
and therefore
$\Pr\bigl[Y\geq x_{S,w}/2\bigr]\;\geq\;\Pr\bigl[Y\geq q_{S,w}/2\bigr]\;\geq\;\frac{1}{2}.$
Whenever $Y\geq x_{S,w}/2$, the unconstrained algorithm sells at least $x_{S,w}/2$ of the $\lfloor x_{S,w}\rfloor$ copies at price $w$. Hence
$\mathbb{E}\left[\textnormal{{UAlg}}\text{ value from }S\right]\;\geq\;\frac{1}{2}\cdot\frac{x_{S,w}}{2}\cdot w\;=\;\frac{1}{4}\,x_{S,w}\,w\;>\;\frac{1}{8}\,\textnormal{{FracOpt}}_{S}.$
Combining all cases, for every desired set $S$ we have
$\mathbb{E}\left[\textnormal{{UAlg}}\text{ value from }S\right]\;\geq\;\frac{1}{8}\,\textnormal{{FracOpt}}_{S}.$
Summing over $S$ and using $\sum_{S}\textnormal{{FracOpt}}_{S}=\textnormal{{FracOpt}}_{\gamma}$ yields
$\mathbb{E}_{M,I}\left[\textnormal{{UAlg}}\right]\;=\;\sum_{S}\mathbb{E}\left[\textnormal{{UAlg}}\text{ value from }S\right]\;\geq\;\frac{1}{8}\,\textnormal{{FracOpt}}_{\gamma},$
which is exactly the statement of Lemma 3.3. ∎
C.2 Proof of Lemma 3.4
Proof.
Fix a desired set $S$ and let $\mathcal{A}=\mathcal{A}_{S}$. We analyze the contribution to the load on a fixed item $e\in S$ coming from buyers in $\mathcal{A}$. As before, we assume that buyers in $\mathcal{A}$ only purchase menu entries corresponding to the bundle $S$; the general case follows from price subadditivity (Appendix B).
For $b\in\mathcal{A}$ let $L_{b}(e)$ be the indicator that buyer $b$ contributes to the load on item $e$, and set $L_{S}(e)=\sum_{b\in\mathcal{A}}L_{b}(e)$. We inspect each of the menu cases of Step 2.
Case 1: $|\mathcal{A}|$ copies at price $w+1$.
Here a buyer accepts if and only if its realized value is $v>w$. Thus
$L_{S}(e)=\sum_{b\in\mathcal{A}}\mathbbm{1}{\left\{v_{b}>w\right\}}.$
Taking expectations,
$\mathbb{E}_{M,I}\left[L_{S}(e)\right]=\sum_{b\in\mathcal{A}}\sum_{v>w}q_{b,v}\leq\sum_{b\in\mathcal{A}}\sum_{v\in\mathcal{S}}x_{b,v}.$
Case 2: one copy at price $w$ w.p. $x_{S,w}/q_{S,w}$, and $|\mathcal{A}|$ copies at $w+1$.
A buyer contributes to $L_{S}(e)$ either by (i) obtaining the (possibly offered) $w$-copy, or (ii) having value $v>w$. Since at most one $w$-copy can be claimed,
$L_{S}(e)\;\leq\;\mathbbm{1}{\left\{\text{$w$-copy is offered}\right\}}\cdot\mathbbm{1}{\left\{\exists b\,:\,v_{b}=w\right\}}\;+\;\sum_{b\in\mathcal{A}}\mathbbm{1}{\left\{v_{b}>w\right\}}.$
Taking expectations,
$\mathbb{E}_{M,I}\left[L_{S}(e)\right]\leq\frac{x_{S,w}}{q_{S,w}}\cdot q_{S,w}+\sum_{b\in\mathcal{A}}\sum_{v>w}q_{b,v}=x_{S,w}+\sum_{b\in\mathcal{A}}\sum_{v>w}q_{b,v}=\sum_{b\in\mathcal{A}}\sum_{v\in\mathcal{S}}x_{b,v}.$
Case 3: one copy at price $w$ w.p. $x_{S,w}$, and $|\mathcal{A}|$ copies at $w+1$.
Exactly as above,
$L_{S}(e)\leq\mathbbm{1}{\left\{\text{$w$-copy offered}\right\}}+\sum_{b\in\mathcal{A}}\mathbbm{1}{\left\{v_{b}>w\right\}},$
and hence
$\mathbb{E}_{M,I}\left[L_{S}(e)\right]\leq x_{S,w}+\sum_{b\in\mathcal{A}}\sum_{v>w}q_{b,v}=\sum_{b\in\mathcal{A}}\sum_{v\in\mathcal{S}}x_{b,v}.$
Case 4: $\lfloor x_{S,w}\rfloor$ copies at price $w$ and $|\mathcal{A}|$ copies at $w+1$.
In this case,
$L_{S}(e)\leq\lfloor x_{S,w}\rfloor+\sum_{b\in\mathcal{A}}\mathbbm{1}{\left\{v_{b}>w\right\}},$
and
$\mathbb{E}_{M,I}\left[L_{S}(e)\right]\leq\lfloor x_{S,w}\rfloor+\sum_{b\in\mathcal{A}}\sum_{v>w}q_{b,v}\leq\sum_{b\in\mathcal{A}}\sum_{v\in\mathcal{S}}x_{b,v}.$
Summing over all desired sets $S$ concludes the proof. Moreover, in each case $L_{S}(e)$ is equal or upper bounded by a sum of negatively associated (NA) $\{0,1\}$ random variables ($Z_{i}$ in the lemma statement). Indeed, for Cases $1,3,4$ the proof is straightforward (note that in these cases the random variables are actually independent).
For Case 2 we will use standard properties from of negative dependence (see Section 2 in [DubhashiRanjan1996]), first set $A_{b}=\mathbbm{1}{\left\{v_{b}=w\right\}}$, $B_{b}=\mathbbm{1}{\left\{v_{b}>w\right\}}$, and $O=\mathbbm{1}{\left\{\text{$w$-copy offered}\right\}}$. Then $A_{b}+B_{b}\leq 1$, so $\{A_{b},B_{b}\}$ is NA. By independent union, $\{A_{b},B_{b}\}_{b\in\mathcal{A}}$ is NA. Applying the non-decreasing map $(A_{b})_{b}\mapsto\mathbbm{1}{\left\{\sum_{b}A_{b}\geq 1\right\}}$ gives $\{\mathbbm{1}{\left\{\exists b:\,v_{b}=w\right\}},\{B_{b}\}_{b}\}$ is NA. Since $O$ is independent, adding $O$ preserves NA. Applying the non-decreasing map $(e,f)\mapsto e\cdot f$ yields
$\Big\{\mathbbm{1}{\left\{\text{$w$-copy offered}\right\}}\cdot\mathbbm{1}{\left\{\exists b:\,v_{b}=w\right\}},\ \{\mathbbm{1}{\left\{v_{b}>w\right\}}\}_{b\in\mathcal{A}}\Big\}\ \text{is NA, done.}$
Finally, $L_{M,I}(e)=\sum_{S}L_{S}(e)$ inherits the same property since random variables for different sets are independent. This enables the concentration bounds we use in the analysis and finishes the proof of the lemma. ∎
Appendix D Graph Setting
D.1 Load Condition (Proof of Lemma 6.1)
Proof.
The proof is similar to the single-minded case. First, fix a desired $(s,t)$ pair and let $\mathcal{A}=\{b:(s_{b},t_{b})=(s,t)\}$. We analyze the contribution to the load on a fixed edge $e$ coming from buyers in $\mathcal{A}$. As before, we assume that buyers in $\mathcal{A}$ only purchase menu entries corresponding to the bundle $(s,t)$; the general case follows from price subadditivity (Appendix D.2).
For $b\in\mathcal{A}$ let $L_{b}(e)$ be the indicator that buyer $b$ contributes to the load on edge $e$, and set $L_{(s,t)}(e)=\sum_{b\in\mathcal{A}}L_{b}(e)$. We also define $\mathcal{P}_{s,t}^{e}=\{\,p\in\mathcal{P}_{s,t}:e\in p\,\}$, the set of $(s,t)$ paths that contain edge $e$. Lastly, let $x_{>v,p}=\sum_{b\in\mathcal{A}}\sum_{v^{\prime}>v}x_{b,v^{\prime},p}$ and define $x_{>v},\,q_{>v}$ similarly. We inspect each of the menu cases of Step 2.
Case 1: $|\mathcal{A}|$ copies at price $w+1$.
Here a buyer accepts if and only if its realized value is $v>w$. Thus
$L_{(s,t)}(e)=\sum_{b\in\mathcal{A}}\mathbbm{1}{\left\{v_{b}>w\right\}}\mathbbm{1}{\left\{\text{$b$ routed through a path in $\mathcal{P}_{s,t}^{e}$}\right\}}.$
Taking expectations,
$\mathbb{E}_{M,I}\left[L_{(s,t)}(e)\right]=q_{>w}\sum_{p\in\mathcal{P}_{s,t}^{e}}\frac{x_{>w,p}}{x_{>w}}=\sum_{p\in\mathcal{P}_{s,t}^{e}}x_{>w,p}\leq\sum_{b\in\mathcal{A}}\sum_{p\in\mathcal{P}_{s,t}^{e}}\sum_{v\in\mathcal{S}}x_{b,v,p}$
Case 2: one copy at price $w$ w.p. $x_{w}/q_{w}$, and $|\mathcal{A}|$ copies at $w+1$.
A buyer contributes to $L_{(s,t)}(e)$ either by (i) obtaining the (possibly offered) $w$-copy, or (ii) having value $v>w$. Since at most one $w$-copy can be claimed,
[…]
$\;+\;\sum_{b\in\mathcal{A}}\mathbbm{1}{\left\{v_{b}>w\right\}}\cdot\mathbbm{1}{\left\{\text{$b$ routed through a path in $\mathcal{P}_{s,t}^{e}$}\right\}}.$
Taking expectations,
$\mathbb{E}_{M,I}\left[L_{(s,t)}(e)\right]$ $\leq\frac{x_{w}}{q_{w}}\cdot q_{w}\cdot\sum_{p\in\mathcal{P}_{s,t}^{e}}\frac{x_{w,p}}{x_{w}}+q_{>w}\sum_{p\in\mathcal{P}_{s,t}^{e}}\frac{x_{>w,p}}{x_{>w}}$
$=\sum_{p\in\mathcal{P}_{s,t}^{e}}x_{w,p}+\sum_{p\in\mathcal{P}_{s,t}^{e}}x_{>w,p}=\sum_{b\in\mathcal{A}}\sum_{p\in\mathcal{P}_{s,t}^{e}}\sum_{v\in\mathcal{S}}x_{b,v,p}$
Case 3: one copy at price $w$ w.p. $x_{w}$, and $|\mathcal{A}|$ copies at $w+1$.
Exactly as above,
$L_{(s,t)}(e)$ $\leq\mathbbm{1}{\left\{\text{$w$-copy offered}\right\}}\cdot\mathbbm{1}{\left\{\text{copy routed through a path in $\mathcal{P}_{s,t}^{e}$}\right\}}$
$+\sum_{b\in\mathcal{A}}\mathbbm{1}{\left\{v_{b}>w\right\}}\cdot\mathbbm{1}{\left\{\text{buyer routed through a path in $\mathcal{P}_{s,t}^{e}$}\right\}},$
and hence taking expectations
$\mathbb{E}_{M,I}\left[L_{(s,t)}(e)\right]$ $\leq x_{w}\cdot\sum_{p\in\mathcal{P}_{s,t}^{e}}\frac{x_{w,p}}{x_{w}}+q_{>w}\sum_{p\in\mathcal{P}_{s,t}^{e}}\frac{x_{>w,p}}{x_{>w}}$
$=\sum_{p\in\mathcal{P}_{s,t}^{e}}x_{w,p}+\sum_{p\in\mathcal{P}_{s,t}^{e}}x_{>w,p}=\sum_{b\in\mathcal{A}}\sum_{p\in\mathcal{P}_{s,t}^{e}}\sum_{v\in\mathcal{S}}x_{b,v,p}$
Case 4: $\lfloor x_{w}\rfloor$ copies at price $w$ and $|\mathcal{A}|$ copies at $w+1$.
In this case,
$L_{(s,t)}(e)$ $\leq\sum_{j=1}^{\left\lfloor x_{w}\right\rfloor}\mathbbm{1}{\left\{\text{$j$-th copy routed through a path in $\mathcal{P}_{s,t}^{e}$}\right\}}$
$+\sum_{b\in\mathcal{A}}\mathbbm{1}{\left\{v_{b}>w\right\}}\cdot\mathbbm{1}{\left\{\text{buyer routed through a path in $\mathcal{P}_{s,t}^{e}$}\right\}},$
and taking expectations
$\mathbb{E}_{M,I}\left[L_{(s,t)}(e)\right]\leq\left\lfloor x_{w}\right\rfloor\cdot\sum_{p\in\mathcal{P}_{s,t}^{e}}\frac{x_{w,p}}{x_{w}}+\sum_{p\in\mathcal{P}_{s,t}^{e}}x_{>w,p}\leq\sum_{b\in\mathcal{A}}\sum_{p\in\mathcal{P}_{s,t}^{e}}\sum_{v\in\mathcal{S}}x_{b,v,p}$
Summing over all desired $(s,t)$ paths concludes the proof. Moreover, in each case $L_{(s,t)}(e)$ is equal or upper bounded by a sum of independent or negatively associated (NA) $\{0,1\}$ random variables. Indeed, for Cases $1,3,4$ the proof is straightforward.
For Case 2, we proceed exactly as in the single-minded case and we get that
$\Big\{\mathbbm{1}{\left\{\text{$w$-copy offered}\right\}}\cdot\mathbbm{1}{\left\{\exists b:\,v_{b}=w\right\}},\ \{\mathbbm{1}{\left\{v_{b}>w\right\}}\}_{b\in\mathcal{A}}\Big\}\ \text{is NA.}$
and since we can multiply each element with independent binary random variables and still preserve NA, we get the result.
Finally, $L_{M,I}(e)=\sum_{(s,t)}L_{(s,t)}(e)$ inherits the same property since random variables for different sets are independent. This enables the concentration bounds we use in the analysis and finishes the proof of the lemma. ∎
D.2 Subadditivity
We now relax the simplifying assumption that a buyer who desires an $(s,t)$-path is only interested on $(s,t)$-path copies. The proof is similar to the single-minded case. Similarly, a buyer could purchase several paths $p_{1},\ldots,p_{\ell}$, that correspond to $(s_{1},t_{1}),\ldots,(s_{\ell},t_{\ell})$ types respectively, from the posted menu such that $\bigcup_{i}p_{i}\supseteq p\in\mathcal{P}_{s,t}$. We will show that after applying essentially the same simple normalization step to the fractional LP solution as in the single-minded case, no buyer can ever strictly prefer such a cover over purchasing an $(s,t)$-path directly. In other words, the posted prices are subadditive.
Pre-processing.
For the graph setting, the objective of the pre-processing step becomes
$\min\;\Bigl(\textsc{FracWeight}(x)+\sum_{e\in E}w_{e}(x)\Bigr)\qquad\text{s.t. }x\in\mathcal{H},$
where
$\textsc{FracWeight}(x)=\sum_{b}\sum_{v\in\mathcal{S}}\sum_{p\in\mathcal{P}_{b}}x_{b,v,p}\quad\text{and}\quad w_{e}(x)=\sum_{b}\sum_{v\in\mathcal{S}}\sum_{p\in\mathcal{P}_{b}\,:\,e\in p}x_{b,v,p}.$
We now state the equivalent of Lemma˜B.1 for the graph setting.
Lemma D.1.
Fix a desired $(s,t)$-path such that $x_{v}<q_{v}$. Suppose there exist different $(s_{i},t_{i})$-paths $p_{1},\dots,p_{\ell}$ such that $\bigcup_{i=1}^{\ell}p_{i}\supseteq p$, for some $p\in\mathcal{P}_{s,t}$. If $x_{(s_{i},t_{i}),v_{i},p_{i}}>0$ for all $i\in[\ell]$, then
$\sum_{i=1}^{\ell}v_{i}>v.$
The proof proceeds exactly as in the single-minded case, except that we transfer mass from paths rather than from sets.
Subadditivity.
We slightly adjust our notation for prices, writing them as $r$ rather than $p$, so as to avoid confusion with paths. As in the single-minded setting, fix a buyer whose desired routing is an $S=(s,t)$-path. Suppose that in the posted menu they purchase paths $p_{1},\dots,p_{\ell}$, that correspond to $T_{i}=(s_{i},t_{i})$ pairs, such that
$\bigcup_{i}p_{i}\supseteq p\quad\text{and}\quad\sum_{i=1}^{\ell}r(T_{i})<r(S)$
for some $p\in\mathcal{P}_{s,t}$. Note that since we always offer $|\mathcal{A}_{s,t}|$ copies of $(s,t)$-paths at some price, a copy of an $(s,t)$-path will always be available as in the single-minded setting.
First, if $r(S)=1$ then $r(T_{i})=0$, contradiction. Now notice that $x_{T_{i},r(T_{i})}>0$ and $x_{S,r(S)-1}$ is not tight by the definition of our pricing rules. Now note that we cannot directly apply Lemma˜D.1 as in the single-minded setting since $x_{T_{i},r(T_{i})}>0$ does not directly imply $x_{T_{i},r(T_{i}),p_{i}}>0$. However, we can overcome this issue via a mass-transfer argument.
Indeed, if $r(T_{i})$ is a crucial value for $T_{i}$ then since path $p_{i}$ is realized we get $x_{T_{i},r(T_{i}),p_{i}}>0$ from the definition of the roulette-drawing, thus we are done in this case. Suppose now that $r(T_{i})$ is tight and $x_{T_{i},r(T_{i}),p_{i}}=0$. Then, since $p_{i}$ is realized it holds that $x_{T_{i},\geq r(T_{i}),p_{i}}>0$ by the definition of the roulette drawing. Thus, there exists a value $v>r(T_{i})$ such that $x_{T_{i},v,p_{i}}>0$. Also, since $r(T_{i})$ is tight there exists a path $p_{i}^{\prime}$ such that $x_{T_{i},r(T_{i}),p_{i}^{\prime}}>0$. We now transfer sufficiently small mass from $x_{T_{i},v,p_{i}}>0$ to $x_{T_{i},r(T_{i}),p_{i}}=0$ and from $x_{T_{i},r(T_{i}),p_{i}^{\prime}}>0$ to $x_{T_{i},v_{i},p_{i}^{\prime}}$. Note that these transfers do not interfere with the feasibility, optimality or the preprocessing step, since they occur between same paths and we balance the value exchanges. Thus, we can assume that $x_{T_{i},r(T_{i}),p_{i}}>0$ for each $i\in[\ell]$.
As a result, we can apply Lemma˜D.1 and get contradiction as in the single-minded case.
Appendix E Lower Bounds
E.1 Proof of Theorem 1.6
Here we show how to remove the $\ln m$ loss in this special case by using Theorem 3 in [poljak]. The theorem states that if $t$ is a prime power and $t^{2}-t$ divides ntn-t, then
$g(n,t,2)\;\geq\;t^{\frac{n-t}{t^{2}-t}}.$
We restrict to the case where the number of items is a cube of a prime, $m=n=p^{3}$ for a sufficiently large prime $p$, and set $t\coloneqq p,\ r\coloneqq 2$. Since $t=p$ is a prime (hence a prime power) and $t^{2}-t=p^{2}-p$ divides $n-t=p^{3}-p$, the condition of Theorem 3 in [poljak] holds, and we obtain
$g(n,t,2)\;\geq\;t^{\frac{n-t}{t^{2}-t}}\;=\;p^{p+1}=t^{t+1}.$
In particular, $g(m,t,2)\;\geq\;t^{t+1}\;\geq\;t^{t}$, thus we select $N=t^{t}$ (the number of partitions in our family).
We now reuse exactly the same construction and analysis as in the proof of Theorem 1.4, specialized to $B=1$ ($r=2$), with this choice of $m$ and $t$. Lemma G.3 gives $\mathbb{P}\!\left[X\geq 1\right]\;\geq\;0.5$, since $Np=B$.
The structural part is exactly as in the $B$-copies proof, with $B=1$ and $r=2$: the analogue of Observation 7.4 (for $B=1$) says that no feasible allocation can serve buyers from more than one group; the analogue of Observation 7.5 (with $B=1$) says that if $E_{i}$ holds, the offline optimum can serve all $t$ buyers in group $i$. Thus, repeating the offline argument from Theorem 1.4 with $B=1$,
$\mathbb{E}\left[\mathrm{OPT}\right]\;\geq\;t\cdot\mathbb{P}\!\left[X\geq 1\right]\;\geq\;0.5\,t.$
Similarly, the same argument from the proof of Theorem 1.4 with $B=1$ shows that each group has total expected value $1$ and, once the algorithm takes value from a group, the expected remaining value in that group is at most 1; combined with the fact that at most one group can be partially served when $B=1$, this yields
$\mathbb{E}\left[\mathrm{ALG}\right]\;\leq\;2.$
Therefore, for $m=p^{3}$ and all sufficiently large primes $p$,
$\frac{\mathbb{E}\left[\mathrm{OPT}\right]}{\mathbb{E}\left[\mathrm{ALG}\right]}\;\geq\;\frac{0.5\,t}{2}=\Omega(t)=\Omega(p)=\Omega(m^{1/3})$
which proves Theorem 1.6.
E.2 Proof of Theorem 1.7
We first note that one can obtain a lower bound of this form by translating the lower-bound instance of Saxena, Velusamy, and Weinberg [SVW22] for prophet inequalities over $q$ partition matroids into a $q$-single-minded instance and counting the number of items.
Instead of going through this translation, we give a direct combinatorial proof, by connecting the Alon–Alweiss measure (AAM) to qualitatively independent partitions and then plugging the resulting partitions into the single-minded lower-bound template used in the proof of Theorem 1.6.
AAM $\Rightarrow$ QI partitions.
Fix a prime number $p$. For an integer $\ell\geq 1$, a set $\mathcal{C}\subseteq\mathbb{Z}_{p}^{\ell}$ is called a $\mathbb{Z}_{p}$-covering $(\ell,p)$-family if for all distinct $u,v\in\mathcal{C}$ and all $s\in\mathbb{Z}_{p}$ there exists an index $i\in[\ell]$ such that
$u_{i}-v_{i}\equiv s\pmod{p}$.
For an integer $N\geq 1$, the Alon–Alweiss measure $\textsc{AAM}(p,N)$ is the smallest $\ell$ for which there exists such a covering family $\mathcal{C}$ with $|\mathcal{C}|\geq N$. Let $g(n,t,2)$ denote the largest size of a family of pairwise (i.e. $2$-way) qualitatively independent $t$-partitions of an $n$-element set.
Lemma E.1.
For every prime $p$ and every integer $N\geq 1$,
$g\bigl(p\cdot\textsc{AAM}(p,N),\,p,\,2\bigr)\;\geq\;N.$
Proof.
Let $\ell:=\textsc{AAM}(p,N)$. By definition there exists a $\mathbb{Z}_{p}$-covering $(\ell,p)$-family $\mathcal{C}\subseteq\mathbb{Z}_{p}^{\ell}$ with $|\mathcal{C}|\geq N$, i.e. for all distinct $u,v\in\mathcal{C}$ and all $s\in\mathbb{Z}_{p}$ there is $i\in[\ell]$ such that $u_{i}-v_{i}\equiv s\pmod{p}$. Set $X:=[\ell]\times\mathbb{Z}_{p}$, so $|X|=p\ell$. For each $u=(u_{1},\dots,u_{\ell})\in\mathcal{C}$ define a $p$-partition $\mathcal{P}_{u}$ of $X$ by blocks
$B_{u}(a)\;:=\;\{(i,a+u_{i})\ :\ i\in[\ell]\}\qquad(a\in\mathbb{Z}_{p}),$
where addition is mod $p$. Then $\{B_{u}(a):a\in\mathbb{Z}_{p}\}$ is a partition of $X$ since every $(i,j)\in X$ lies in the unique block with $a\equiv j-u_{i}\pmod{p}$.
We claim that $\mathcal{P}_{u}$ and $\mathcal{P}_{v}$ are qualitatively independent for all distinct $u,v\in\mathcal{C}$. Fix $a,b\in\mathbb{Z}_{p}$ and put $s:=b-a\in\mathbb{Z}_{p}$. By the covering property, there exists $i\in[\ell]$ such that $u_{i}-v_{i}\equiv s\pmod{p}$. Then
$(i,a+u_{i})\in B_{u}(a)\quad\text{and}\quad a+u_{i}\equiv a+v_{i}+s\equiv b+v_{i}\pmod{p}$,
hence $(i,a+u_{i})=(i,b+v_{i})\in B_{v}(b)$ and so $B_{u}(a)\cap B_{v}(b)\neq\emptyset$. Since $a,b$ were arbitrary, $\mathcal{P}_{u}$ and $\mathcal{P}_{v}$ are qualitatively independent. Thus $\{\mathcal{P}_{u}:u\in\mathcal{C}\}$ is a family of $|\mathcal{C}|\geq N$ pairwise qualitatively independent $p$-partitions of an $n=p\ell=p\cdot\textsc{AAM}(p,N)$ element set. Therefore $g\bigl(p\cdot\textsc{AAM}(p,N),\,p,\,2\bigr)\geq N$. ∎
Saxena, Velusamy, and Weinberg [SVW22] show that there exists an absolute constant $\alpha>0$ such that for all sufficiently large primes $p$,
$\textsc{AAM}(p,p^{p})\;\leq\;p^{\,2-\alpha/\log\log p}.$
Applying Lemma E.1 with $N=p^{p}$ and $t:=p$, we obtain
$g\bigl(t\cdot\textsc{AAM}(t,t^{t}),\,t,\,2\bigr)\;\geq\;t^{t}$
and, writing $m:=t\cdot\textsc{AAM}(t,t^{t})$, we have
$m\;\leq\;t^{\,3-\alpha/\log\log t}.$
This implies that for all sufficiently large such $m$ there exists an absolute constant $c>0$ with
$t\;\geq\;m^{\,1/3+c/\log\log m}.$ (1)
Lower Bound Instance.
For such an $m$ and $t$ we now take a ground set of $m$ items and fix a family
$\mathcal{F}=\bigl\{\mathcal{P}^{(1)},\dots,\mathcal{P}^{(N)}\bigr\},\qquad N=t^{t},$
of $N$ pairwise qualitatively independent $t$-partitions of the items, guaranteed by Lemma E.1.
We then instantiate exactly the single-minded lower-bound construction from the proof of Theorem 1.6 for $B=1$ using this family $\mathcal{F}$. The structural and probabilistic analysis of this instance (number of "fully active" groups, offline prophet value, and the bound on any online algorithm) is identical to the $B=1$ case of Theorem 1.6 and implies:
$\mathbb{E}[\mathrm{OPT}]\geq 0.5t\qquad\text{and}\qquad\mathbb{E}[\mathrm{ALG}]\leq 2,$
for every online algorithm. Thus
$\frac{\mathbb{E}[\mathrm{OPT}]}{\mathbb{E}[\mathrm{ALG}]}=0.25t\;=\;\Omega\!\left(m^{1/3+c/\log\log m}\right)$
by (1), for all sufficiently large $m$ of the form $m=t\cdot\textsc{AAM}(t,t^{t})$ with $t$ prime.
Appendix F Tree Setting
In this section, we show how to achieve an $O(1)$ competitive ratio for the case where single-minded buyers demand paths in unit-capacitated trees and have $\{0,1\}$ valuations. While this result is already known due to Chawla et al. [chawla18] (see their more general Theorem 1.4), we include it here to illustrate how our framework can recover existing results, thereby demonstrating its versatility and power.
Formally, consider a tree $T=(V,E)$ with unit capacities on all edges and set an arbitrary node $r\in V$ to be its root. Each buyer’s value is realized according to a Bernoulli distribution, i.e., $\mathcal{D}_{b}=\mathrm{Be}(q_{b})$. Using the same techniques as in [chawla18], one could partition buyers’ values into $\log(v_{\max}/v_{\min}^{*})$ valuation classes to achieve their $O(\log H)$ competitive ratio. For clarity, however, we do not perform this step here, since it is straightforward and already established; we focus on the analysis as a proof of concept to illustrate our framework.
Approach.
The approach follows our general algorithmic framework and the analysis borrows the main ideas from [chawla18]. First, solve the ex-ante linear relaxation with appropriately scaled capacities (Step 1). For trees, we select $\gamma=10$. Next, post the randomized menu resulting from Step 2. We will now prove that the above (randomized) menu obtains a constant competitive ratio.
Definitions.
We say that a buyer $b$ is blocked by another buyer $b^{\prime}\neq b$ if $S_{b}$ and $S_{b^{\prime}}$ share an edge and $b^{\prime}$ is accepted by the constrained algorithm before $b$ arrives. Let $\textnormal{{Alg}}_{M,I}$ be the set of buyers accepted by the constrained algorithm, and define $\textsc{Blocked}_{M,I}\;=\;\textnormal{{UAlg}}_{M,I}\setminus\textnormal{{Alg}}_{M,I},$ the set of buyers accepted by the unconstrained algorithm but rejected by the constrained one due to blocking.
For a buyer $b$, let $u_{b}$ be the node of $S_{b}$ that is closest to the root $r$. Any edge of $S_{b}$ incident to $u_{b}$ is called a peak edge of $b$; the remaining edges of $S_{b}$ are non-peak. The parent of an edge is the unique adjacent edge closer to the root. Let $e_{b}^{1}$ and $e_{b}^{2}$ denote the peak edges of $b$ (one or two, depending on the structure of $S_{b}$), and define $f_{b}$ as the unique edge that is the parent of $e_{b}^{1}$; if $e_{b}^{2}$ also exists, then this edge is also the parent of $e_{b}^{2}$. We partition blocked buyers according to where the blocking occurs:
  • $\textsc{Peak}_{M,I}$ is the set of buyers in $\textsc{Blocked}_{M,I}$ that are blocked on one of their peak edges,
  • $\textsc{NonPeak}_{M,I}$ is the set of buyers in $\textsc{Blocked}_{M,I}$ that are blocked only on non-peak edges.
Finally, let $\ell_{A}(e)$ denote the load of edge $e$ due to buyers in $A$. Note that $\ell_{\textnormal{{UAlg}}_{M,I}}(e)=L_{M,I}(e)$.
Analysis.
Combining Lemma˜3.4 and Lemma˜3.3 with our choice of $\gamma=10$ (and using the fact that $c(e)=1$ for all edges), we obtain the following.
Corollary F.1.
For every edge $e\in E$,
$\mathbb{E}_{M,I}\left[L_{M,I}(e)\right]\;\leq\;\frac{c(e)}{10}\;=\;0.1.$
Moreover,
$\mathbb{E}_{M,I}\left[\textnormal{{UAlg}}_{M,I}\right]\;\geq\;\frac{\textnormal{{FracOpt}}_{\gamma}}{8}\;\geq\;\frac{\textnormal{{FracOpt}}}{80}.$
For a fixed $M$ and $I$ it holds that (note that buyers are $\{0,1\}$ thus only the cardinality of the sets matters):
${}\textnormal{{Alg}}_{M,I}=\textnormal{{UAlg}}_{M,I}-\textsc{Blocked}_{M,I}=\textnormal{{UAlg}}_{M,I}-\textsc{Peak}_{M,I}-\textsc{NonPeak}_{M,I}$ (2)
We first prove the following lemma.
Lemma F.2.
For every edge $e$ and buyer $b$ it holds that:
$\mathbb{E}_{M,I}\left[\ell_{\textnormal{{UAlg}}_{M,I}\setminus\{b\}}(e)\mid b\in\textnormal{{UAlg}}_{M,I}\right]\leq\mathbb{E}_{M,I}\left[\ell_{\textnormal{{UAlg}}_{M,I}}(e)\right]$
Proof.
Fix an edge $e$, a buyer $b$ and the set $\mathcal{A}$ of buyers that desire the same set of edges, $S_{b}$. Define also $\mathcal{A}_{M,I}\coloneqq\textnormal{{UAlg}}_{M,I}\cap\mathcal{A}$. It suffices to prove that:
$\mathbb{E}_{M,I}\left[\ell_{\mathcal{A}_{M,I}\setminus\{b\}}(e)\mid b\in\textnormal{{UAlg}}_{M,I}\right]\leq\mathbb{E}_{M,I}\left[\ell_{\mathcal{A}_{M,I}}(e)\right]$
since whether $b\in\textnormal{{UAlg}}_{M,I}$ does not affect buyers that desire sets different from $S_{b}$ (as we will formalize when we establish subadditivity of the menu).
First, if we offer $|\mathcal{A}|$ copies of bundle $S_{b}$ then for every buyer $b^{\prime}\in\mathcal{A}$ the event $b^{\prime}\in\mathcal{A}_{M,I}$ depends only on whether buyer $b^{\prime}$ is realized thus the lemma trivially holds. Otherwise, one copy of bundle $S_{b}$ is offered with some probability. Since $b\in\mathcal{A}_{M,I}$, buyer $b$ got the single copy of bundle $S_{b}$, thus $\ell_{\mathcal{A}_{M,I}\setminus\{b\}}(e)=0$, proving the lemma. ∎
We now prove the following two lemmas.
Lemma F.3.
$\mathbb{E}_{M,I}\left[\textsc{NonPeak}_{M,I}\right]\leq 0.1\,\mathbb{E}_{M,I}\left[\textnormal{{UAlg}}_{M,I}\right]$
Proof.
It holds that:
[…]
$\leq\sum_{b\in\textnormal{{UAlg}}_{M,I}}\ell_{\textnormal{{UAlg}}_{M,I}\setminus\{b\}}(f_{b})$
$\leq\sum_{b\in\mathcal{N}}\mathbbm{1}{\left\{b\in\textnormal{{UAlg}}_{M,I}\right\}}\ell_{\textnormal{{UAlg}}_{M,I}\setminus\{b\}}(f_{b})$
Taking expectations in both parts we get:
$\mathbb{E}_{M,I}\left[\textsc{NonPeak}_{M,I}\right]$ $\leq\sum_{b\in\mathcal{N}}\mathbb{P}\!\left[b\in\textnormal{{UAlg}}_{M,I}\right]\mathbb{E}_{M,I}\left[\ell_{\textnormal{{UAlg}}_{M,I}\setminus\{b\}}(f_{b})\mid b\in\textnormal{{UAlg}}_{M,I}\right]$
$\leq\sum_{b\in\mathcal{N}}\mathbb{P}\!\left[b\in\textnormal{{UAlg}}_{M,I}\right]\mathbb{E}_{M,I}\left[\ell_{\textnormal{{UAlg}}_{M,I}}(f_{b})\right]$
$=0.1\cdot\mathbb{E}_{M,I}\left[\textnormal{{UAlg}}_{M,I}\right],$
where in the second inequality we used Lemma˜F.2. ∎
Lemma F.4.
$\mathbb{E}_{M,I}\left[\textsc{Peak}_{M,I}\right]\leq 0.2\;\mathbb{E}_{M,I}\left[\textnormal{{UAlg}}_{M,I}\right]$
Proof.
It holds that:
[…]
$\leq\sum_{b\in\textnormal{{UAlg}}_{M,I}}\mathbbm{1}{\left\{\ell_{\textnormal{{Alg}}_{M,I}}(e_{b}^{1})>0\right\}}+\mathbbm{1}{\left\{\ell_{\textnormal{{Alg}}_{M,I}}(e_{b}^{2})>0\right\}}$
$\leq\sum_{b\in\mathcal{N}}\mathbbm{1}{\left\{b\in\textnormal{{UAlg}}_{M,I}\right\}}\left(\ell_{\textnormal{{UAlg}}_{M,I}\setminus\{b\}}(e_{b}^{1})+\ell_{\textnormal{{UAlg}}_{M,I}\setminus\{b\}}(e_{b}^{2})\right)$
Taking expectations in both parts we get similarly to the non-peak case:
$\mathbb{E}_{M,I}\left[\textsc{Peak}_{M,I}\right]$ $\leq\sum_{b\in\mathcal{N}}\mathbb{P}_{M,I}\!\left[b\in\textnormal{{UAlg}}_{M,I}\right]\left(\mathbb{E}_{M,I}\left[\ell_{\textnormal{{UAlg}}_{M,I}}(e_{b}^{1})\right]+\mathbb{E}_{M,I}\left[\ell_{\textnormal{{UAlg}}_{M,I}}(e_{b}^{2})\right]\right)$
$\leq\sum_{b\in\mathcal{N}}\mathbb{P}_{M,I}\!\left[b\in\textnormal{{UAlg}}_{M,I}\right]\left(0.1+0.1\right)$
$=0.2\;\mathbb{E}_{M,I}\left[\textnormal{{UAlg}}_{M,I}\right]$
From Corollary˜F.1, Lemma˜F.4, Lemma˜F.3 and taking expectations in Equation˜2 we get:
$\mathbb{E}_{M,I}\left[\textnormal{{Alg}}_{M,I}\right]\geq 0.7\,\mathbb{E}_{M,I}\left[\textnormal{{UAlg}}_{M,I}\right]\geq\frac{7}{800}\textnormal{{FracOpt}}$
Appendix G Deferred Proofs
G.1 Proof of Lemma 2.1
Proof.
Consider the optimal allocation for a fixed realization of buyers’ values. For each buyer $b$ and value $v\in\mathcal{S}$, let $y_{b,v}\in\{0,1\}$ indicate whether in this realization buyer $b$ is allocated bundle $S_{b}$ when their value is $v$. By feasibility of the offline solution, for every $e\in\mathcal{M}$ we have
$\sum_{b:e\in S_{b}}\sum_{v\in\mathcal{S}}y_{b,v}\leq c(e),$
and moreover $y_{b,v}=0$ unless buyer $b$ realizes value $v$, implying $\mathbb{E}_{I}\left[y_{b,v}\right]\leq q_{b,v}$, where the expectation is over the randomness in the buyers’ realized values. Define the fractional allocation $x_{b,v}:=\mathbb{E}_{I}\left[y_{b,v}\right]$. By construction, $0\leq x_{b,v}\leq q_{b,v}$, so the box constraints of the LP are satisfied. Taking expectations of the offline capacity constraints yields, for every $e\in\mathcal{M}$,
$\sum_{b:e\in S_{b}}\sum_{v\in\mathcal{S}}x_{b,v}=\mathbb{E}_{I}\left[\sum_{b:e\in S_{b}}\sum_{v\in\mathcal{S}}y_{b,v}\right]\leq c(e),$
so the LP capacity constraints are also satisfied. The objective value of this feasible fractional solution is
$\sum_{b\in[n]}\sum_{v\in\mathcal{S}}x_{b,v}\cdot v=\mathbb{E}_{I}\left[\sum_{b\in[n]}\sum_{v\in\mathcal{S}}y_{b,v}\cdot v\right]=\textnormal{{Opt}},$
where the last equality holds because the offline allocation attains the optimal welfare for each realization. Since FracOpt is the optimal value of the LP and we constructed a feasible fractional solution with value Opt, we obtain $\textnormal{{FracOpt}}\geq\textnormal{{Opt}}$. ∎
G.2 Proof of Lemma 2.3
Proof.
Fix a buyer $b$ and consider the vector $X^{(b)}=(x_{b,v})_{v\in\mathcal{S}}$, where values in $\mathcal{S}$ are ordered increasingly. Recall the value‐exchange argument from the main text: whenever $v^{\prime}>v$, we may shift an infinitesimal amount of mass from $x_{b,v}$ to $x_{b,v^{\prime}}$ without violating any LP constraints, unless one of the box constraints $0\leq x_{b,v}$ or $x_{b,v^{\prime}}\leq q_{b,v^{\prime}}$ becomes tight. Such a shift strictly increases the objective because $v^{\prime}>v$.
Case 1: A crucial value $w$ exists.
Suppose there exists a value $v<w$ with $x_{b,v}>0$. Since $v<w$ and $x_{b,w}<q_{b,w}$, we may shift mass from $x_{b,v}$ to $x_{b,w}$ without violating feasibility, strictly improving the objective - a contradiction. Thus $x_{b,v}=0$ for all $v<w$. Next, suppose there exists a value $u>w$ with $x_{b,u}<q_{b,u}$. Since $x_{b,w}>0$, we may shift mass from $x_{b,w}$ to $x_{b,u}$, again increasing the objective. Hence all $u>w$ must satisfy $x_{b,u}=q_{b,u}$.
Case 2: No crucial value exists.
Here every $v$ satisfies either $x_{b,v}=0$ or $x_{b,v}=q_{b,v}$. Suppose there exist $u>v$ with $x_{b,u}=0$ and $x_{b,v}=q_{b,v}$. Since $u>v$ and $x_{b,u}$ is not tight, shifting mass from $v$ to $u$ preserves feasibility and increases the objective - a contradiction. Thus all zero entries occur at smaller values, and all tight entries occur at larger values, with no interleaving. ∎
G.3 Proof of Lemma 2.4
Proof.
Fix a bundle $S$ and let $\mathcal{A}=\mathcal{A}_{S}$ and $X=X^{(S)}$ for simplicity. Recall that
$x_{S,v}\;=\;\sum_{b\in\mathcal{A}}x_{b,v}\qquad\text{and}\qquad q_{S,v}\;=\;\sum_{b\in\mathcal{A}}q_{b,v},$
and we refer to $x_{S,v}$ as tight if $x_{S,v}=q_{S,v}$ and crucial if $0<x_{S,v}<q_{S,v}$. We begin with two structural claims.
Lemma G.1.
If any buyer $b\in\mathcal{A}$ has a crucial value, then all crucial values across buyers in $\mathcal{A}$ occur at the same value. We call this value the crucial value of $X$.
Proof.
Suppose buyer $b$ has crucial value $w_{b}$ and buyer $b^{\prime}$ has crucial value $w_{b^{\prime}}$ with $w_{b}<w_{b^{\prime}}$. Because both buyers demand the same bundle $S$, shifting a small amount of mass from $x_{b,w_{b}}$ to $x_{b^{\prime},w_{b^{\prime}}}$ preserves all LP capacity constraints: both variables consume exactly the same set of items. This shift increases the objective because $w_{b^{\prime}}>w_{b}$, contradicting optimality of the LP solution. Hence no two buyers in $\mathcal{A}$ can have crucial values at different levels, proving the claim. ∎
Lemma G.2.
If $X$ has a crucial value $w$, then
$x_{S,v}=0\;\text{ for all }v<w,\qquad\text{and}\qquad x_{S,v}=q_{S,v}\;\text{ for all }v>w.$
Proof.
Since $w$ is crucial for $X$, we have $0<x_{S,w}<q_{S,w}$. First suppose there exists $v<w$ with $x_{S,v}>0$. Then there is some buyer $b\in\mathcal{A}$ with $x_{b,v}>0$. Because $x_{S,w}<q_{S,w}$, there exists some buyer $b^{\prime}\in\mathcal{A}$ with $x_{b^{\prime},w}<q_{b^{\prime},w}$.
By the same value–exchange argument used in Lemma 2.3, we can shift an infinitesimal amount of mass from $x_{b,v}$ to $x_{b^{\prime},w}$: this preserves all LP constraints, since both variables correspond to the same bundle $S$ and all item capacities are affected in exactly the same way. The objective strictly increases because $w>v$, contradicting optimality. Hence $x_{S,v}=0$ for all $v<w$.
Now suppose there exists $u>w$ with $x_{S,u}<q_{S,u}$. Then there is a buyer $b^{\prime}\in\mathcal{A}$ with $x_{b^{\prime},u}<q_{b^{\prime},u}$. Since $x_{S,w}>0$, there exists a buyer $b\in\mathcal{A}$ with $x_{b,w}>0$. Again, applying the value–exchange argument, we can shift an infinitesimal amount of mass from $x_{b,w}$ to $x_{b^{\prime},u}$. Feasibility is preserved (both entries use the same bundle $S$), and the objective strictly increases because $u>w$, contradicting optimality. Thus no such $u$ can exist, and we must have $x_{S,u}=q_{S,u}$ for all $u>w$. This proves the claim. ∎
We now conclude the proof of Lemma˜2.4. If no crucial value exists in $X$, then for every $v$ we have either $x_{S,v}=0$ or $x_{S,v}=q_{S,v}$, and by the same exchange argument as above, zeros must occur only at lower values and tight values only at higher values. If a crucial value exists, then by Lemma˜G.1 it is unique, and by Lemma˜G.2 $X$ consists of zeros below it, the crucial value, and tight entries above it. Thus in all cases the vector $X$ has exactly the form stated in the lemma. ∎
G.4 Proof of Lemma 7.3
Proof.
Since $B<\ln m$ and $m\geq 16$, we have
$\left(\frac{m}{2B\ln m}\right)^{1/(B+2)}>1,$
and therefore
$1\;\leq\;t\;\leq\;\left(\frac{m}{2B\ln m}\right)^{1/(B+2)}\;\leq\;m^{1/(B+2)}\;\leq\;m.$
Thus $t$ is a well-defined integer in $[m]$. Applying Theorem 7.2 with $n=m$ and $r=B+1$ gives
$g(m,t,B+1)\;\geq\;\frac{e(B+1)}{t}\exp\!\left(\frac{m}{(B+1)t^{B+1}}\right).$
By the definition of $t$,
$t^{B+2}\;\leq\;\frac{m}{2B\ln m}.$
therefore
$\frac{m}{(B+1)t^{B+1}}\;\geq\;\frac{m}{(B+1)\,\frac{m}{2B\ln m}\,\frac{1}{t}}\;=\;\frac{2Bt\ln m}{B+1}.$
Plugging this into the bound on $g(m,t,B+1)$ yields
$g(m,t,B+1)\;\geq\;\frac{e(B+1)}{t}\exp\!\left(\frac{2Bt\ln m}{B+1}\right)\;=\;\frac{e(B+1)}{t}\,m^{2Bt/(B+1)}.$
It remains to show that
$\frac{e(B+1)}{t}\,m^{2Bt/(B+1)}\;\geq\;Bt^{t}.$
Taking logarithms and using $t\leq m^{1/(B+2)}$ (and $m$ sufficiently large) shows that the left-hand side grows much faster than the right-hand side as a function of $m$. Hence the inequality holds, and
$g(m,t,B+1)\;\geq\;Bt^{t},$
as claimed. ∎
G.5 Proof of Lemma 7.7
Proof.
Apply Lemma˜7.6 with $n=m=dt$, $r=B+1$, and $d=n/t$:
$g_{\mathrm{bal}}(m,t,B+1,d)\;\geq\;\frac{e(B+1)}{t}\cdot\exp\!\left(\frac{m}{(B+1)t^{B+1}}\right)=\frac{e(B+1)}{t}\cdot\exp\!\left(\frac{d}{(B+1)t^{B}}\right).$
By the definition of $t$,
$t^{B+1}\;\leq\;\frac{d}{2B\ln d}\quad\Rightarrow\quad\frac{d}{(B+1)t^{B}}\;\geq\;\Omega(t\ln d),$
Hence
$g_{\mathrm{bal}}(m,t,B+1,d)\;\geq\;d^{\Omega(t)}.$
Since $t\leq d$, we have $d^{\Omega(t)}\geq Bt^{t}$ for all sufficiently large $d$, proving the claim. ∎
G.6 Auxiliary Lemma
Lemma G.3.
For a random variable $X$ with mean $\mu$ that is the sum of some independent random variables on $\{0,1\}$ we get:
$\mathbb{P}\!\left[X\geq\mu/2\right]\geq 0.5\cdot\min\{1,\mu\}$
Proof.
Let $X=\sum_{i=1}^{n}X_{i}$, where the $X_{i}$ are independent $\{0,1\}$-valued random variables with $\mathbb{P}\!\left[X_{i}=1\right]=p_{i}$. Then $\mu=\mathbb{E}\left[X\right]=\sum_{i=1}^{n}p_{i}$.
Case 1: $0<\mu\leq 1$.
Here $\mu/2\in(0,1/2]$, hence (since $X$ is integer-valued) $\{X\geq\mu/2\}=\{X\geq 1\}$. Thus
$\mathbb{P}\!\left[X\geq\mu/2\right]=\mathbb{P}\!\left[X\geq 1\right]=1-\mathbb{P}\!\left[X=0\right]=1-\prod_{i=1}^{n}(1-p_{i})\geq 1-\exp\left(-\sum_{i=1}^{n}p_{i}\right)=1-e^{-\mu}.$
using $1-x\leq e^{-x}$ for $x\in[0,1]$. For $\mu\in[0,1]$, it holds that $1-e^{-\mu}\geq\mu/2$ therefore we get the lemma.
Case 2: $\mu\geq 1$.
It is a standard fact for Poisson binomial random variables (i.e., sums of independent Bernoulli variables) that any median $m$ of $X$ satisfies $\lfloor\mu\rfloor\leq m\leq\lceil\mu\rceil$ (see for example Theorem 3.2 and Corollary 3.1 in [jogdeo]). In particular, $m\geq\lfloor\mu\rfloor$. Since $\mu\geq 1$ implies $\lfloor\mu\rfloor\geq\mu/2$, we obtain
$\mathbb{P}\!\left[X\geq\mu/2\right]\;\geq\;\mathbb{P}\!\left[X\geq m\right]\;\geq\;1/2.$
Combining the two cases, we conclude that
$\mathbb{P}\!\left[X\geq\mu/2\right]\geq 0.5\cdot\min\{1,\mu\}.$